### v2

# Global well-posedness for small data in a 3D temperature-velocity model with Dirichlet boundary noise

Abstract.
We study a three-dimensional Boussinesq-type temperature-velocity system on a bounded smooth domain $\mathcal{D}\subset\mathbb{R}^{3}$, where the velocity $u^{\varepsilon}$ solves the Navier-Stokes equations and the temperature $\theta^{\varepsilon}$ is driven by Dirichlet boundary noise of intensity $\sqrt{\varepsilon}$. The boundary forcing produces a stochastic convolution $Z^{\varepsilon}$ which is, in general, only continuous in time with values in $H^{-\frac{1}{2}-\delta_{\theta}}(\mathcal{D})$. To handle this roughness together with initial data $\theta_{0}\in W^{s,6/5}(\mathcal{D})$, we work in the ambient space $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$ with $\delta_{u}\geq\max\{\delta_{\theta},\frac{1}{2}-s\}$.
Given a finite time $T>0$, for any $p>4$ and sufficiently small initial data, we prove existence and uniqueness of a mild solution $(u^{\varepsilon},\theta^{\varepsilon})$ up to a stopping time $\tau^{\varepsilon}\leq T$ such that
$u^{\varepsilon}\in W^{1,p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))\cap L^{p}(0,\tau^{\varepsilon};H^{\frac{3}{2}-\delta_{u}}(\mathcal{D})),\quad\theta^{\varepsilon}\in C(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})).$
Moreover, we obtain a high-probability global existence estimate of the form $\mathbb{P}(\tau^{\varepsilon}=T)\geq 1-C\varepsilon$, with $C=C(\delta_{\theta},T)>0.$
Key words and phrases:
Navier-Stokes Equations, Stochastic Boundary Conditions
1991 Mathematics Subject Classification:
60H15, 60H30, 76D03
1. Introduction
Let $\mathcal{D}\subset\mathbb{R}^{3}$ be a bounded, open domain (i.e. non-empty, connected set) with smooth boundary $\partial\mathcal{D}$, and let $T>0$ be a fixed final time. We study the coupled velocity-temperature system for an incompressible fluid in $\mathcal{D}$, modelled by Navier-Stokes equations with thermal advection-diffusion, and perturbed by Dirichlet boundary noise. Concretely, for each $\varepsilon>0$ we consider
[…]
where $u^{\varepsilon}\colon\mathcal{D}\times(0,T)\to\mathbb{R}^{3}$ is the fluid velocity, $\theta^{\varepsilon}\colon\mathcal{D}\times(0,T)\to\mathbb{R}$ is the temperature field, $p^{\varepsilon}\colon\mathcal{D}\times(0,T)\to\mathbb{R}$ is the pressure, $e_{3}=(0,0,1)^{T}$, and $(W_{t})_{t\geq 0}$ is a $\mathcal{Q}$-Wiener process acting only on the boundary with intensity which scales depending on the parameter $\varepsilon$.
The parameter $\varepsilon$ influences the probability of global existence up to time $T$. Indeed, the three-dimensional (3D) Navier-Stokes equations are only known to be globally well‑posed under smallness conditions on initial data and forcing (see, e.g., [8, Chapter 9]). Since stochastic boundary forcing enters the velocity equation through buoyancy coupling, we impose a small‑noise regime in order to retain a nonzero probability of well‑posedness on $[0,T]$. However, even for arbitrarily small $\varepsilon>0$, the boundary noise may assume large values with positive probability. Consequently, we construct the solution $(u^{\varepsilon},\theta^{\varepsilon})$ up to a random stopping time $\tau^{\varepsilon}\leq T$, defined to interrupt the evolution if the stochastic forcing becomes too large.
A key difficulty of the model comes from the fact that Dirichlet boundary noise is much rougher than the forcing acting in the interior of the domain. Even in the linear heat equation with boundary noise, the stochastic convolution produced by the boundary forcing typically lives only slightly below the threshold $H^{-1/2}(\mathcal{D})$. We quantify this loss through a small exponent $\delta_{\theta}>0$, which measures how far the noise-driven component falls into $H^{-1/2-\delta_{\theta}}(\mathcal{D})$.
To couple this temperature field back into the three-dimensional Navier-Stokes equation via the buoyancy term $\theta^{\varepsilon}e_{3}$, we solve the fluid problem in a (slightly) weaker Sobolev scale $H^{-1/2-\delta_{u}}(\mathcal{D})$, chosen so that both the noise contribution and the initial temperature can be interpreted as forcing terms at the fluid level. This leads to the compatibility requirement
$\delta_{u}\geq\max\{\delta_{\theta},\ \tfrac{1}{2}-s\},$
where $s\in[0,1/2)$ is the Sobolev regularity of the initial temperature $\theta_{0}\in W^{s,6/5}(\mathcal{D})$.
Under this condition (and for $p>4$ in the maximal-regularity framework), we prove the existence and uniqueness of the coupled system on the interval $0,\tau^{\varepsilon}]$, satisfying
$u^{\varepsilon}\in W^{1,p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))\cap L^{p}(0,\tau^{\varepsilon};H^{\frac{3}{2}-\delta_{u}}(\mathcal{D})),\quad\theta^{\varepsilon}\in C(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})).$
Furthermore, we establish the high-probability estimate $\mathbb{P}(\tau^{\varepsilon}=T)\geq 1-C\varepsilon$. As $s\uparrow 1/2$, the constraint from the initial datum disappears and one may choose $\delta_{u}=\delta_{\theta}$, recovering a symmetric regularity regime.
It is worth underlining two aspects of the boundary noise: its mathematical difficulties, and its physical meaning. Firstly, Da Prato and Zabczyk were the first to rigorously demonstrate in [9] that solutions to the heat equation subject to Dirichlet boundary noise exhibit significantly reduced regularity compared to solutions driven by interior stochastic forcing, even in one spatial dimension. Specifically, due to the presence of boundary noise, the best regularity that one can hope for, even in the linear case, is continuity in time with values in a negative-order Bessel potential space, namely $C_{t}(H_{x}^{-\frac{1}{2}-\delta_{\theta}})$ for any $\delta_{\theta}>0$. This limitation reflects the difficulties of noisy Dirichlet boundary conditions.
The low spatial regularity of the stochastic convolution in the Dirichlet setting is a major obstruction for non-linear problems: even basic products and non-linear maps may fail to be well-defined in a classical sense, and standard energy methods or fixed-point arguments often require additional structure. As a consequence, the available well-posedness theory for non-linear PDEs with Dirichlet white-noise boundary data is rather limited and typically concerns only a limited class of non-linearities; see for instance [13, 23, 26] and references therein. Other works treat situations where the boundary forcing is regularised (e.g. coloured in space and/or fractional in time) or otherwise smoothed at the boundary, which can lead to better regularity and allow one to handle genuinely non-linear dynamics; see [1, 14]. In our setting the boundary noise enters the temperature equation linearly, but the resulting rough temperature acts as a forcing in the Navier-Stokes component, and its interplay with the convective non-linearity is one of the main analytic difficulties. It is also worth noting that the $H^{-\frac{1}{2}-\delta_{\theta}}$ limitation is closely related to a boundary-layer singularity: solutions to the associated linear Dirichlet problem may blow up as one approaches $\partial\mathcal{D}$, while remaining smooth in the interior, see [3, 4, 7, 16].
In parallel, the velocity field $u^{\varepsilon}$ enjoys, in this context, the maximal regularity of the Stokes operator in the same low-regularity framework introduced above. This constitutes the optimal regularity according to the maximal regularity theory for the Stokes operator on $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$, with suitable divergence-free and boundary conditions on the spaces. Our maximal regularity approach follows the framework developed by Prüss and Wilke in [25, Section 5].
Secondly, from a physical point of view, stochastic boundary forcing encapsulates the influence of unresolved, rapid fluctuations, such as boundary-layer instabilities, or small-scale convection, that cannot be parametrized deterministically at the scale of our model. This philosophy traces back to Hasselmann’s stochastic climate paradigm [18], in which slow, large-scale dynamics are driven by fast, random perturbations. The boundary-noise framework thus provides a mathematically tractable way to “close” the system. In this spirit, we plan to investigate in a future work in which sense the system (1.1) arises as a limit of a multiscale fast-slow model.
Lastly, a natural direction for future research is the analysis of the two-dimensional counterpart of the system. In the 2D setting, it is an interesting open problem to determine whether global well-posedness can be established on any time interval without relying on the stopping time $\tau^{\varepsilon}$ (i.e., independent of the noise intensity).
2. Preliminaries and main results
In this section we introduce the notation which will be used throughout the paper, describe our approach to solve the coupled velocity-temperature problem (1.1), and present the main results of our work.
2.1. Notations and functional setting
We work on a complete filtered probability space $\left(\Omega,\mathcal{F},\left(\mathcal{F}_{t}\right)_{t},\mathbb{P}\right)$. A stochastic process $\Phi$, taking values in a measurable space, is adapted if $\Phi_{t}$ is $\mathcal{F}_{t}$-measurable for any $t\geq 0$. It is progressively measurable if the map $(s,\omega)\mapsto\Phi_{s}(\omega)$ is measurable on $([0,t]\times\Omega,\mathcal{B}(0,t)\otimes\mathcal{F}_{t})$ for every $t\geq 0$, with $\mathcal{B}(0,t)$ being the Borel $\sigma$-algebra on $[0,t].$
Let $\mathcal{D}\subset\mathbb{R}^{3}$ be a bounded, open domain (i.e. non-empty, connected set) with smooth boundary. We denote by $(W_{t})_{t}$ a $\mathcal{Q}$-Wiener process on $L^{2}(\partial\mathcal{D})$ defined on $\left(\Omega,\mathcal{F},\left(\mathcal{F}_{t}\right)_{t},\mathbb{P}\right)$, and represented by
$W(t,x)=\sum_{k\geq 0}\lambda_{k}e_{k}(x)\beta_{k}(t),\quad t\geq 0,\ x\in\partial\mathcal{D},$
where $(e_{k})_{k}\subset L^{2}(\partial\mathcal{D})$ is an orthonormal basis of $L^{2}(\partial\mathcal{D})$, $(\beta_{k}(t))_{k}$ are independent Brownian motions, and $(\lambda_{k})_{k}$ are the non-negative square roots of the eigenvalues of the covariance operator $\mathcal{Q}$. For further details, we refer to [11].
To handle the stochastic forcing on the boundary, we employ the Dirichlet map $D$, following the approach of [9]. It is defined as the linear map
$D\colon L^{2}(\partial\mathcal{D})\to L^{2}(\mathcal{D}),\quad Dh:=u,$
where $u$ denotes the unique weak solution to the Dirichlet problem
$\left\{\begin{aligned} \Delta u&=0,\quad&x\in&\mathcal{D},\\ u|_{\partial\mathcal{D}}&=h,\quad\quad&x\in&\partial\mathcal{D}.\end{aligned}\right$.
Regarding the functional setting, for $s\in\mathbb{R}$ and $p\in(1,\infty),$ we denote by $W^{s,p}(\mathcal{D})$ the fractional Sobolev space and by $H^{s,p}(\mathcal{D})$ the Bessel potential space. In particular, we set $H^{0,p}(\mathcal{D})=L^{p}(\mathcal{D})$, and write $H^{s}(\mathcal{D})=H^{s,2}(\mathcal{D})$. Note that $H^{s,p}(\mathcal{D})=W^{s,p}(\mathcal{D})$ for any $s\in\mathbb{N}.$
We denote by
$\Delta\colon H^{2,2}(\mathcal{D})\cap H^{1,2}_{0}(\mathcal{D})\subset L^{2}(\mathcal{D})\to L^{2}(\mathcal{D})$
the Dirichlet Laplacian, and we consider the fractional powers $(-\Delta)^{\alpha}$, for any $\alpha\in(0,1)$. Their domains are given by
$D((-\Delta)^{\alpha}):=\prescript{}{0}{H}^{2\alpha}(\mathcal{D}):=\left\{\begin{aligned} &H^{2\alpha}(\mathcal{D}),&\quad&0<\alpha<\frac{1}{4},\\ &H^{\frac{1}{2}}_{00}(\mathcal{D}),&\quad&\alpha=\frac{1}{4},\\ &H^{2\alpha}_{0}(\mathcal{D}),&\quad&\frac{1}{4}<\alpha\leq\frac{1}{2},\\ &H^{2\alpha}(\mathcal{D})\cap H^{1}_{0}(\mathcal{D}),&\quad&\frac{1}{2}<\alpha<1,\end{aligned}\right$.
where $H^{2\alpha}_{00}(\mathcal{D})$ consists of all $u\in H^{\frac{1}{2}}(\mathcal{D})$ such that
[…]
with $\rho(x)$ being any $C^{\infty}$ function comparable to $dist(x,\partial\mathcal{D})$; see [20, 28] for more details on the domain of the fractional powers of the Dirichlet Laplacian.
We define the space of solenoidal (divergence-free) square-integrable vector fields by
$L^{2}_{\sigma}(\mathcal{D})=\overline{\left\{u\in C^{\infty}_{c}(\mathcal{D};\mathbb{R}^{3})\ :\ \mathrm{div}(u)=0\text{ in }\mathcal{D}\right\}}^{\norm{\cdot}_{L^{2}(\mathcal{D})}}.$
The space $L^{2}(\mathcal{D})$ can be decomposed as
$L^{2}(\mathcal{D})=L^{2}_{\sigma}(\mathcal{D})\oplus G_{2}(\mathcal{D}),$
where $G_{2}(\mathcal{D}):=\left\{u\in L^{2}(\mathcal{D})\ :\ u=\nabla\pi,\text{ for some }\pi\in H^{1,2}_{loc}(\mathcal{D})\right\}$, and there exists a unique projection $P\colon L^{2}(\mathcal{D})\to L^{2}_{\sigma}(\mathcal{D})$ called the Helmholtz projection in $L^{2}(\mathcal{D})$. We denote by $A$ the Stokes operator with Dirichlet boundary conditions, defined by
$Au:=-P\Delta u,\quad D(A):=H^{2,2}(\mathcal{D})\cap H^{1,2}_{0}(\mathcal{D})\cap L^{2}_{\sigma}(\mathcal{D}).$
In the Hilbert space $X_{0}=L^{2}_{\sigma}(\mathcal{D})$, the Stokes operator $A$ is a non-negative, self-adjoint operator with compact inverse; see, for instance, [27, Section 2]. In particular, $A$ is sectorial of angle 0 and A-A generates a bounded analytic $C_{0}$-semigroup on $X_{0}$; see, for example, [12, 19]. Moreover, since $A$ is a non-negative self-adjoint operator on a Hilbert space, it admits a bounded $\mathcal{H}^{\infty}$-functional calculus of angle 0; see [17, Chapters 2 and 7]. In the specific case of the Stokes operator, the domains of the fractional powers $A^{\alpha}$ in $L^{2}_{\sigma}(\mathcal{D})$ have been identified in [15] and are given by
$D(A^{\alpha})=\prescript{}{0}{H^{2\alpha}}(\mathcal{D})\cap L^{2}_{\sigma}(\mathcal{D}),\qquad\alpha\in(0,1).$
For notational convenience, we introduce the following spaces:
$H^{s}_{\sigma}(\mathcal{D})=H^{s}(\mathcal{D})\cap L_{\sigma}^{2}(\mathcal{D}),\quad\prescript{}{0}{H}^{s}_{\sigma}(\mathcal{D})=\prescript{}{0}{H^{s}(\mathcal{D})}\cap L^{2}_{\sigma}(\mathcal{D}),\quad s\geq 0,$
and, by duality
$H^{-s}_{\sigma}(\mathcal{D})=\left(H^{s}(\mathcal{D})\cap L^{2}_{\sigma}(\mathcal{D})\right)^{\prime},\quad\prescript{}{0}{H}^{-s}_{\sigma}(\mathcal{D})=\left(\prescript{}{0}{H^{s}(\mathcal{D})}\cap L^{2}_{\sigma}(\mathcal{D})\right)^{\prime},\quad s>0.$
We will work with an extension of the classical Stokes operator to a weaker setting, which we denote by $A_{w}$. Specifically, for any sufficiently small $\delta_{u}>0$, we define the weak Stokes operator by
(2.1) $A_{w}\colon\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})$ $\to\prescript{}{0}{H^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})}$
$\langle A_{w}u,v\rangle$ $=\langle\nabla u,\nabla v\rangle_{{H^{\frac{1}{2}-\delta_{u}}},{H^{-\frac{1}{2}+\delta_{u}}}},$
for all pairs $(u,v)\in\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\times\prescript{}{0}{H^{\frac{1}{2}+\delta_{u}}_{\sigma}(\mathcal{D})}$. Here, $\langle\cdot,\cdot\rangle_{{H^{\frac{1}{2}-\delta_{u}}},{H^{-\frac{1}{2}+\delta_{u}}}}$ denotes the duality pairing between $H^{\frac{1}{2}-\delta_{u}}(\mathcal{D})$} and $H^{-\frac{1}{2}+\delta_{u}}(\mathcal{D})$}.
Lastly, given a final time $T>0$ and $p\in(1,\infty)$, we set
$\mathbb{E}_{t,p}^{\delta_{u}}:=W^{1,p}(0,t;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))\cap L^{p}(0,t;\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})),$
for any $t\in[0,T]$. This space will be the maximal regularity space on which the velocity of our model evolves. Additionally, we define the real interpolation space
$V_{p}^{\delta_{u}}=\left(\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}),\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\right)_{1-\frac{1}{p},p}.$
For more information on these function spaces, we refer the reader to [20, 28, 21].
2.2. Description of our approach
The approach we adopt to study model (1.1) is the following. First, we decouple the model (1.1) into a system for temperature only, given a prescribed fluid velocity; and a system for velocity only, given a prescribed temperature. Namely, we study first the equation for the temperature $\theta^{\varepsilon}\colon\mathcal{D}\times(0,T)\to\mathbb{R}$, which reads as
(2.2) $\left\{\begin{aligned} \partial_{t}\theta^{\varepsilon}+u\cdot\nabla\theta^{\varepsilon}-\Delta\theta^{\varepsilon}&=0,\quad&\text{ in }&\mathcal{D}\times(0,T),\\ \theta^{\varepsilon}|_{\partial\mathcal{D}}&=\sqrt{\varepsilon}\frac{dW}{dt},\quad&\text{ in }&\partial\mathcal{D}\times(0,T),\\ \theta^{\varepsilon}|_{t=0}&=\theta_{0},\quad&\text{ in }&\mathcal{D}.\end{aligned}\right$.
Here, $u\colon\mathcal{D}\times(0,T)\to\mathbb{R}^{3}$ is a given velocity and $\theta_{0}$ is the initial condition. Second, we study the 3D Navier-Stokes equations for the velocity $u\colon\mathcal{D}\times(0,T)\to\mathbb{R}^{3}$ given by
[…]
Here, $f^{\theta}\colon\mathcal{D}\times(0,T)\to\mathbb{R}^{3}$ is a prescribed forcing, $p\colon\mathcal{D}\times(0,T)\to\mathbb{R}$ is the pressure, and $u_{0}$ denotes the initial condition.
Then, following the idea in [10, 2], we split the analysis of (2.2) into two problems, corresponding to the decomposition $\theta_{t}^{\varepsilon}=Z_{t}^{\varepsilon}+\zeta_{t}^{\varepsilon}.$ The first term $Z^{\varepsilon}_{t}$ corresponds to the stochastic linear problem with non-homogeneous Dirichlet boundary conditions given by
(2.4) $\left\{\begin{aligned} \partial_{t}Z^{\varepsilon}&=\Delta Z^{\varepsilon},\quad&\text{ in }&\mathcal{D}\times(0,T),\\ Z^{\varepsilon}|_{\partial\mathcal{D}}&=\sqrt{\varepsilon}\frac{dW}{dt},\quad&\text{ in }&\partial\mathcal{D}\times(0,T),\\ Z^{\varepsilon}|_{t=0}&=0,\quad&\text{ in }&\mathcal{D}.\end{aligned}\right$.
The solution of the previous equation can be interpreted in mild form, as in [9, 16]. Next, we consider the remainder $\zeta_{t}^{\varepsilon}:=\theta_{t}^{\varepsilon}-Z_{t}^{\varepsilon}$, which satisfies
(2.5) $\left\{\begin{aligned} \partial_{t}\zeta^{\varepsilon}+u\cdot\nabla\zeta^{\varepsilon}+u\cdot\nabla Z^{\varepsilon}-\Delta\zeta^{\varepsilon}&=0,\quad&\text{ in }&\mathcal{D}\times(0,T),\\ \zeta^{\varepsilon}|_{\partial\mathcal{D}}&=0,\quad&\text{ in }&\partial\mathcal{D}\times(0,T),\\ \zeta^{\varepsilon}|_{t=0}&=\theta_{0},\quad&\text{ in }&\mathcal{D}.\end{aligned}\right$.
Here, we interpret again the solution $\zeta^{\varepsilon}\colon\mathcal{D}\times(0,T)\to\mathbb{R}$ in the mild form, namely
$\zeta_{t}^{\varepsilon}=e^{t\Delta}\theta_{0}-\int_{0}^{t}e^{\left(t-r\right)\Delta}\left(u_{r}\cdot\nabla\zeta_{r}^{\varepsilon}\right)dr-\int_{0}^{t}e^{\left(t-r\right)\Delta}\left(u_{r}\cdot\nabla Z_{r}^{\varepsilon}\right)dr.$
We conclude this section by describing the solution of the linear temperature problem associated to (2.2), i.e. by recalling the following result concerning the mild solution of (2.4), which follows from [9] and the factorisation trick for the stochastic convolution [11, Section 5.3.1].
Proposition 2.1.
For any $\delta_{\theta}>0$, fix
$\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2},\quad\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4}.$
Assume that
(2.6) $\sum_{k}\lambda_{k}^{2}\norm{(-\Delta)^{\beta_{\theta}}De_{k}}_{2}^{2}<\infty.$
The following holds.
  1. (i)
    The process $Z^{\varepsilon}_{t}=-\sqrt{\varepsilon}\int_{0}^{t}\Delta e^{(t-r)\Delta}DdW_{r}$ is the unique $\left(\mathcal{F}_{t}\right)_{t}$-adapted mild solution of (2.4), and takes values in $H^{-2\alpha_{\theta}}(\mathcal{D})$.
  2. (ii)
    The process
    (2.7) $\xi^{\varepsilon}_{t}:=(-\Delta)^{-\alpha_{\theta}}Z^{\varepsilon}_{t}=\sqrt{\varepsilon}\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}}e^{(t-r)\Delta}DdW_{r},$
    has a continuous version in $L^{2}(\mathcal{D})$ and $Z^{\varepsilon}$ has a continuous version in $H^{-2\alpha_{\theta}}(\mathcal{D})$.
  3. (iii)
    Let $T>0$. For any $\vartheta>0$, there exists a positive constant $C=C(\delta_{\theta},T)$ such that
    $\mathbb{P}\left(\sup_{0\leq t\leq T}\norm{Z_{t}^{\varepsilon}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}>\vartheta\right)\leq\frac{\varepsilon}{\vartheta^{2}}C\sum_{k}\lambda_{k}^{2}\norm{(-\Delta)^{\beta_{\theta}}De_{k}}_{2}^{2}.$
We remark that point (iii) of the previous result is key since it controls, depending on the noise intensity $\sqrt{\varepsilon}$, the probability of the sup norm of $Z_{t}^{\varepsilon}$ to be small in an appropriate Bessel potential space.
2.3. Main results
We are now in a position to state the main results of our work. The first one concerns the well-posedness of the system (2.5).
Theorem 2.2.
Let $s\in[0,1/2)$ and $p>2$. Then there exist $\overline{\delta}_{\theta}=\overline{\delta}_{\theta}(s,p)>0$ and $\overline{\delta}_{u}=\overline{\delta}_{u}(s,p)>0$ such that for every $0<\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $0<\delta_{u}<\overline{\delta}_{u}$, the following holds. Let
$\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2},\quad\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4},$
and denote by $Z^{\varepsilon}_{t}=-\sqrt{\varepsilon}\int_{0}^{t}\Delta e^{(t-r)\Delta}D\,dW_{r}$ the continuous version in $H^{-2\alpha_{\theta}}(\mathcal{D})$ considered in Proposition 2.1. Then, for every divergence–free velocity field
$u\in L^{p}\bigl(0,T;\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\bigr)\cap L^{\infty}\bigl(0,T;\prescript{}{0}{H}^{\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\bigr)$
and every initial datum $\theta_{0}\in W^{s,6/5}(\mathcal{D})$, there exists a unique mild solution $\zeta^{\varepsilon}$ of (2.5) of the form
$\zeta_{t}^{\varepsilon}=e^{t\Delta}\theta_{0}-\int_{0}^{t}e^{(t-r)\Delta}\bigl(u_{r}\cdot\nabla\zeta_{r}^{\varepsilon}\bigr)\,dr-\int_{0}^{t}e^{(t-r)\Delta}\bigl(u_{r}\cdot\nabla Z_{r}^{\varepsilon}\bigr)\,dr,$
with trajectories in
$\zeta^{\varepsilon}\in C(0,T;W^{s,6/5}(\mathcal{D}))\quad\mathbb{P}-\text{a.s.}$
A straightforward but important consequence of the proof of the previous result is the following estimate.
Corollary 2.3.
Under the assumptions of Theorem 2.2, let $\zeta\in C(0,T;W^{s,\frac{6}{5}}(\mathcal{D}))$ and define
$\tilde{\zeta}_{t}:=e^{t\Delta}\theta_{0}-\int_{0}^{t}e^{\left(t-r\right)\Delta}\left(u_{r}\cdot\nabla\zeta_{r}\right)dr-\int_{0}^{t}e^{\left(t-r\right)\Delta}\left(u_{r}\cdot\nabla Z_{r}^{\varepsilon}\right)dr,$
Then, for any $0<\lambda<1-\delta_{u}$, it holds
(2.8) $\begin{split}\norm{\tilde{\zeta}}_{C(0,T;W^{s,\frac{6}{5}}(\mathcal{D}))}&\leq\norm{e^{t\Delta}\theta_{0}}_{C(0,T;W^{s,\frac{6}{5}}(\mathcal{D}))}\\ &\quad+C\norm{u}_{L^{\frac{p}{\lambda+\delta_{u}}}(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))}\left(\norm{\zeta}_{C(0,T;W^{s,\frac{6}{5}}(\mathcal{D}))}+\norm{\xi^{\varepsilon}}_{C(0,T;L^{2}(\mathcal{D}))}\right),\end{split}$
where $C=C(p,\lambda,\delta_{u},\delta_{\theta},s,T)>0$.
So far, in Proposition 2.1, Theorem 2.2, and Corollary 2.3, we have investigated the well-posedness of the temperature equation (2.2). Moreover, we can control the norm of its solution $\theta^{\varepsilon}=Z^{\varepsilon}+\zeta^{\varepsilon}$ with high probability depending on $\varepsilon$, since we are able to estimate the norms of $Z^{\varepsilon}$ (by Proposition 2.1) and $\zeta^{\varepsilon}$ (by Corollary 2.3).
We now turn to the analysis of the 3D Navier-Stokes equations (2.3), assuming that $f^{\theta}$ is a prescribed forcing term. Global well-posedness results for the 3D Navier–Stokes equations are well known under smallness assumptions on the initial datum and the forcing; see for instance [8, Chapter 9]. In the classical energy setting one typically assumes $f^{\theta}\in L^{2}(0,T;L^{2}(\mathcal{D}))$ and $u_{0}\in H^{1}(\mathcal{D})$.
In our coupled model the forcing is of buoyancy type, $f^{\theta}=-\theta^{\varepsilon}e_{3}$, and the temperature is decomposed as $\theta^{\varepsilon}=Z^{\varepsilon}+\zeta^{\varepsilon}$. The noise component $Z^{\varepsilon}$ is, in general, only continuous in time with values in $H^{-\frac{1}{2}-\delta_{\theta}}(\mathcal{D})$, for any $\delta_{\theta}>0$ compatible with the covariance summability condition (2.6); the remainder $\zeta^{\varepsilon}$ is controlled in $W^{s,6/5}(\mathcal{D})$. We therefore work in the ambient space $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$ for the Stokes/Navier-Stokes maximal regularity theory, choosing $\delta_{u}$ so that
$\delta_{u}\geq\max\{\delta_{\theta},\ \tfrac{1}{2}-s\},$
which ensures that both $Z^{\varepsilon}$ and $\zeta^{\varepsilon}$ can be interpreted as forcing terms in $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$. Accordingly, we formulate the Navier-Stokes well-posedness result for general forcing
$f^{\theta}\in L^{p}(0,T;H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})),\quad p>\frac{2}{1-\delta_{u}},$
which is the natural class arising from the maximal regularity framework in the low-regularity setting.
The notion of solution considered for the 3D Navier-Stokes problem (2.3) is the one of mild solution in the weak setting determined by the weak Stokes operator $A_{w}$. We now formalise this notion.
Definition 2.4.
A function $u$ is a solution of the 3D Navier-Stokes equations (2.3) on the time interval $[0,T]$ if
(2.9) $u_{t}=e^{-tA_{w}}u_{0}+\int_{0}^{t}e^{-(t-r)A_{w}}P\left(-u_{r}\cdot\nabla u_{r}+f_{r}^{\theta}\right)dr,$
for any $0\leq t\leq T$, and possesses the regularity
$u\in\mathbb{E}_{T,p}^{\delta_{u}}=W^{1,p}(0,T;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))\cap L^{p}(0,T;\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})).$
Here, and throughout the rest of the paper, $P$ denotes the Helmholtz projection, defined by interpolation from $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$ to $H^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})$. The global existence results for the velocity problem read as follows.
Theorem 2.5 (Global well-posedness for small data).
[…]
$\max\left(\norm{u_{0}}_{V_{p}^{\delta_{u}}},\norm{f^{\theta}}_{L^{p}(0,T;H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}\right)\leq\frac{\eta}{4M},$
then there exists a unique solution $u\in\mathbb{E}_{T,p}^{\delta_{u}}$ with $\|u\|_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq\eta$ of the 3D Navier-Stokes equation (2.3) in the sense of Definition 2.4.
Remark 2.6.
Note that Definition 2.4 is well-posed due to the fact that, as will be demonstrated in Section 4.1, the operator $A_{w}$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})$. Thus, the operator $A_{w}$ generates an analytic semigroup on $\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})$, ensuring that the variation of constants formula (2.9) is meaningful, thanks also to the estimate of the convective term
$\norm{P(u\cdot\nabla u)}_{L^{p}(0,T;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))}\lesssim\norm{u}^{2}_{\mathbb{E}_{T,p}^{\delta_{u}}},$
which will be proved in Lemma 4.4.
We now shift our focus to the fully coupled velocity-temperature system (1.1). The corresponding notion of solution is introduced below.
Definition 2.7.
The triple $(u^{\varepsilon},\theta^{\varepsilon},\tau^{\varepsilon})$ is a solution of (1.1) on $[0,T]$ if $\tau^{\varepsilon}\colon\Omega\to[0,T]$ is a stopping time and $(u^{\varepsilon},\theta^{\varepsilon})$ is a stochastic process with trajectories in
$\mathbb{E}^{\delta_{u}}_{\tau^{\varepsilon},p}\times C\bigl(0,\tau^{\varepsilon};H^{-\tfrac{1}{2}-\delta_{u}}(\mathcal{D})\bigr),$
where $u^{\varepsilon}$ is adapted to $(\mathcal{F}_{t})_{t}$ as a process in $\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})$, and progressively measurable as a process in $\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})$, while $\theta^{\varepsilon}$ is adapted as a process in $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$, and, for any $0\leq t\leq\tau^{\varepsilon}$, it holds
$u_{t}^{\varepsilon}$ $=e^{-tA_{w}}u_{0}+\int_{0}^{t}e^{-(t-r)A_{w}}P\left(-u_{r}^{\varepsilon}\cdot\nabla u_{r}^{\varepsilon}-\theta^{\varepsilon}_{r}e_{3}\right)dr,$
$\theta_{t}^{\varepsilon}$ $=\zeta_{t}^{\varepsilon}+Z_{t}^{\varepsilon},$
$\zeta_{t}^{\varepsilon}$ $=e^{t\Delta}\theta_{0}-\int_{0}^{t}e^{\left(t-r\right)\Delta}\left(u_{r}^{\varepsilon}\cdot\nabla\zeta_{r}^{\varepsilon}\right)dr-\int_{0}^{t}e^{\left(t-r\right)\Delta}\left(u_{r}^{\varepsilon}\cdot\nabla Z_{r}^{\varepsilon}\right)dr,$
$Z_{t}^{\varepsilon}$ $=-\sqrt{\varepsilon}\int_{0}^{t}\Delta e^{(t-r)\Delta}DdW_{r},$
with probability one.
Lastly, we state the main result of our work.
Theorem 2.8.
Let $T>0$ and $s\in[0,\tfrac{1}{2})$. Fix $p>4.$ Choose a parameter $\gamma>\frac{1}{4}$ sufficiently close to $\frac{1}{4}$ such that
$a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}<1,\qquad\lambda_{\max}:=p(1-a(s,\gamma))>\lambda_{\min}:=\frac{1}{2}-s.$
(Such a $\gamma$ always exists since $p>4$ implies p(1a(s,1/4))>1/2sp(1-a(s,1/4))>1/2-s). Choose
$\lambda\in(\lambda_{\min},\,\lambda_{\max})\cap(0,1),$
and define the thresholds
$\overline{\delta}_{\theta}:=\min\{\lambda,\ 2\gamma-\tfrac{1}{2}\},\quad\overline{\delta}_{u}:=\min\{\lambda_{\max}-\lambda,1-\lambda\}.$
Let parameters $\delta_{\theta},\delta_{u}$ satisfy
$0<\delta_{\theta}\leq\overline{\delta}_{\theta},\qquad 0<\delta_{u}<\overline{\delta}_{u},$
subject to the compatibility conditions
$\delta_{u}\geq\max\left\{\delta_{\theta},\ \tfrac{1}{2}-s\right\},\quad\delta_{u}<1-\frac{2}{p}.$
Set
$\alpha_{\theta}:=\tfrac{1}{4}+\tfrac{\delta_{\theta}}{2},\qquad\beta_{\theta}:=\tfrac{1}{4}-\tfrac{\delta_{\theta}}{4},$
and assume
$\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{L^{2}(\mathcal{D})}^{2}<\infty.$
Then there exist constants $\eta>0$ and $\widetilde{M}=\widetilde{M}(\delta_{\theta},\delta_{u},T)\geq 2$ such that the following holds: if
$(\theta_{0},u_{0})\in W^{s,6/5}(\mathcal{D})\times V_{p}^{\delta_{u}},\qquad\max\{\|\theta_{0}\|_{W^{s,6/5}(\mathcal{D})},\ \|u_{0}\|_{V_{p}^{\delta_{u}}}\}\leq\frac{\eta}{16\widetilde{M}}$,
then for every $\varepsilon>0$ there exists a unique solution $(u^{\varepsilon},\theta^{\varepsilon},\tau^{\varepsilon})$ of (1.1) on $[0,T]$ in the sense of Definition 2.7, such that
$\|u^{\varepsilon}\|_{\mathbb{E}^{\delta_{u}}_{\tau^{\varepsilon},p}}\leq\eta,\qquad\|\theta^{\varepsilon}\|_{C\bigl(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})\bigr)}\leq\eta,\qquad\mathbb{P}-\text{a.s.}$
Moreover, there exists a constant $C(\delta_{\theta},T)>0$ such that
$\mathbb{P}(\tau^{\varepsilon}=T)\geq 1-\frac{64\,\widetilde{M}^{2}\,\varepsilon}{\eta^{2}}\,C(\delta_{\theta},T)\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}.$
Remark 2.9 (Consistency of parameters).
Let us briefly verify that the set of admissible parameters is non-empty under the condition $p>4$.
  1. (1)
    Existence of $\lambda$. As $\gamma\downarrow\frac{1}{4}$, we have $a(s,\gamma)\to\frac{3}{4}+\frac{s}{2}$. Consequently, $\lambda_{\max}\to p(\frac{1}{4}-\frac{s}{2})$. The condition $\lambda_{\max}>\lambda_{\min}=\frac{1}{2}-s$ is equivalent to $p>\frac{4(1/2-s)}{1-2s}=2$. Since $p>4$, the interval for $\lambda$ is non-empty for $\gamma$ sufficiently close to $1/4$.
  2. (2)
    Existence of $\delta_{u}$. We need to ensure that the lower bound for $\delta_{u}$ is strictly smaller than the upper bound. Taking $\delta_{\theta}$ sufficiently small, the lower bound is $\frac{1}{2}-s$. On the other hand, the upper bound is $\min\{\lambda_{\max}-\lambda,1-\lambda,1-\frac{2}{p}\}$. We thus need to check that there exists a $\lambda$ such that
    $\frac{1}{2}-s<\min\left\{1-\frac{2}{p},\ \lambda_{\max}-\lambda,\ 1-\lambda\right\}.$
    • First, the condition $\frac{1}{2}-s<1-\frac{2}{p}$ is equivalent to $p>\frac{4}{1+2s}$. Since $s\in[0,1/2)$, the term $\frac{4}{1+2s}$ is at most $4$. Thus, the hypothesis $p>4$ guarantees this condition holds.
    • Second, we check that $\frac{1}{2}-s<\lambda_{\max}-\lambda$. This is equivalent to finding $\lambda$ such that $\lambda<\lambda_{\max}-(\frac{1}{2}-s)$. Since we can choose $\lambda$ arbitrarily close to its lower bound $\lambda_{\min}=\frac{1}{2}-s$, such a $\lambda$ exists if $\lambda_{\min}<\lambda_{\max}-\lambda_{\min}$, i.e., $2\lambda_{\min}<\lambda_{\max}$. Since $\lambda_{\max}=p\bigl(1-a(s,\gamma)\bigr)$ and $a(s,\gamma)\to\frac{3}{4}+\frac{s}{2}$ as $\gamma\downarrow\frac{1}{4}$, we have $\lim\limits_{\gamma\downarrow\frac{1}{4}}\lambda_{\max}=\frac{p}{4}(1-2s);$ note that $\frac{p}{4}(1-2s)>2\lambda_{\min}=1-2s$, thanks to the hypothesis $p>4$. Thus the original condition is verified.
    • Finally, we check that $\frac{1}{2}-s<1-\lambda$. This is equivalent to finding $\lambda$ such that $\lambda<\frac{1}{2}+s$. Since $\lambda$ can be chosen close to $\lambda_{\min}=\frac{1}{2}-s$, i.e. $\lambda=\lambda_{\min}+\varepsilon_{\lambda},$ with $\varepsilon_{\lambda}>0$. Thus the condition can be rewritten as $\frac{1}{2}-s<\frac{1}{2}+s+\varepsilon_{\lambda}$, which holds for any $s\geq 0$.
2.4. Overview
The rest of this work is organised as follows. In Section 3, we present the analysis of the temperature problem (2.2), assuming a given velocity field with specified regularity. The strategy is to split the problem into a linear system with non-homogeneous Dirichlet boundary noise (2.4), and the remainder system (2.5). In Section 4, we analyse the velocity system (2.3), subject to a prescribed temperature forcing of specified regularity. In particular, we show that the weak Stokes operator admits a bounded $\mathcal{H}^{\infty}$-calculus and enjoys maximal regularity. We then apply these results to prove a global well-posedness result for small data for the 3D Navier–Stokes equations (2.3). Lastly, Section 5 is entirely devoted to the proof of Theorem 2.8.
3. Temperature problem
In this section we describe how to solve the temperature problem (2.2), which is split in the linear problem (2.4), and the remainder problem (2.5).
3.1. Stochastic linear problem with non-homogeneous Dirichlet boundary conditions
We start this section by proving the first part of Proposition 2.1, in particular the well-posedness and the regularity for the mild solution $Z_{t}^{\varepsilon}$ of the linear problem (2.4) subject to Dirichlet noise boundary conditions. Note that this result is well-known in the literature, see for instance [9, Proposition 3.1].
Proof of Proposition 2.1 (i)-(ii).
Since (i)-(ii) do not depend on $\varepsilon$, we assume $\varepsilon=1$ and omit the dependence on $\varepsilon$ of $\xi$ and $Z$ in the following.
(i) First, considering the splitting
$Z_{t}=(-\Delta)^{\alpha_{\theta}}\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{(t-r)\Delta}(-\Delta)^{\beta_{\theta}}DdW_{r},$
$Z_{t}$ is well-defined and takes values in $H^{-2\alpha_{\theta}}(\mathcal{D})$, assuming (2.6), if
(3.1) $\left\{\begin{aligned} \alpha_{\theta}+\beta_{\theta}&>\frac{1}{2},\\ \beta_{\theta}&<\frac{1}{4}.\end{aligned}\right$.
Indeed, the first inequality in (3.1) is a consequence of the square integrability conditions needed to define the Itô integral. The second guarantees that
$\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{(t-r)\Delta}(-\Delta)^{\beta_{\theta}}DdW_{r}=\sum_{k}\lambda_{k}\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{(t-r)\Delta}(-\Delta)^{\beta_{\theta}}De_{k}\,d\beta_{k}(r)$
is well-defined. In particular $De_{k}\in D((-\Delta)^{\beta_{\theta}})$, where the explicit characterisation of the domain for fractional powers of the Dirichlet Laplacian was recalled in Section 2.1. Note that (3.1) is satisfied by our choice of $\alpha_{\theta},\beta_{\theta}.$
(ii) We observe that $\xi_{t}$ is a Gaussian process with values in $L^{2}(\mathcal{D})$ and with zero mean. If we are able to prove that there exist $M>0$ and $\eta\in(0,1]$ such that
(3.2) $\mathbb{E}\left[\norm{\xi_{t_{2}}-\xi_{t_{1}}}_{2}^{2}\right]\leq M(t_{2}-t_{1})^{\eta},\quad\forall\,0\leq t_{1}\leq t_{2},$
then the thesis follows from the Kolmogorov test for Gaussian processes, see [11, Proposition 3.16]. Consider
$\quad\xi_{t}=\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}}e^{(t-r)\Delta}DdW_{r}=\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{(t-r)\Delta}(-\Delta)^{\beta_{\theta}}DdW_{r},$
with
$\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2},\quad\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4}.$
Then, for any $0\leq t_{1}\leq t_{2}$, we have
$\xi_{t_{2}}-\xi_{t_{1}}=I_{1}+I_{2},$
with
$I_{1}:=\sum_{k}\lambda_{k}\int_{t_{1}}^{t_{2}}\left[(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{\left(t_{2}-r\right)\Delta}(-\Delta)^{\beta_{\theta}}De_{k}\right]\,d\beta_{k}(r)$
and
$I_{2}:=\sum_{k}\lambda_{k}\int_{0}^{t_{1}}\left[(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}\left(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta}\right)(-\Delta)^{\beta_{\theta}}De_{k}\right]d\beta_{k}(r).$
For the estimate of the first integral $I_{1}$, by the independence of $(\beta_{k})_{k}$, the Itô-isometry, and the estimate for the heat-semigroup, we get
(3.3) $\mathbb{E}\left[\|I_{1}\|_{2}^{2}\right]$ $\leq\sum_{k}\lambda_{k}^{2}\int_{t_{1}}^{t_{2}}\norm{(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{\left(t_{2}-r\right)\Delta}(-\Delta)^{\beta_{\theta}}De_{k}}_{2}^{2}dr$
$\leq C\sum_{k}\lambda_{k}^{2}\norm{(-\Delta)^{\beta_{\theta}}De_{k}}_{2}^{2}\int_{t_{1}}^{t_{2}}(t_{2}-r)^{-2(1-\alpha_{\theta}-\beta_{\theta})}dr.$
By our choice of $\alpha_{\theta}$ and $\beta_{\theta}$, it holds
$2(1-\alpha_{\theta}-\beta_{\theta})=-1+\frac{\delta_{\theta}}{2}.$
Since
$\int_{t_{1}}^{t_{2}}(t_{2}-r)^{-1+\frac{\delta_{\theta}}{2}}\,dr=\frac{2}{\delta_{\theta}}(t_{2}-t_{1})^{{\delta_{\theta}}/2},$
we get, up to renaming $C$,
(3.4) $\begin{split}\mathbb{E}\left[\|I_{1}\|_{2}^{2}\right]\leq C(t_{2}-t_{1})^{\frac{\delta_{\theta}}{2}}\sum_{k}\lambda_{k}^{2}\norm{(-\Delta)^{\beta_{\theta}}De_{k}}_{2}^{2}.\end{split}$
For the estimate of the second integral $I_{2}$, we use the fact that for any $\gamma\in(0,\alpha_{\theta}+\beta_{\theta})$ there exists $C>0$ such that for any $0\leq r\leq t_{1}$, it holds
(3.5) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2}(\mathcal{D}))}\leq C(t_{2}-t_{1})^{\gamma}(t_{1}-r)^{-(1-\alpha_{\theta}-\beta_{\theta}+\gamma)},$
where $\mathcal{L}(L^{2})$ is the space of bounded, linear operators from $L^{2}(\mathcal{D})$ to itself. Indeed, by the semigroup property
$e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta}=e^{(t_{1}-r)\Delta}(e^{(t_{2}-t_{1})\Delta}-I).$
Further, by the commuting property of the fractional powers with the heat semigroup
$(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})=(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I).$
Considering the norms on the previous identity, it yields
(3.6) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2})}\leq\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|_{\mathcal{L}(L^{2})}\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}.$
The first term on the right-hand-side can be estimated by the analyticity of the semigroup as
(3.7) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|_{\mathcal{L}(L^{2})}\leq C(t_{1}-r)^{-1+\alpha_{\theta}+\beta_{\theta}-\gamma}.$
For the second term on the right-hand side, again thanks to the analyticity of the semigroup, it holds
$(e^{(t_{2}-t_{1})\Delta}-I)g=\Delta\int_{0}^{t_{2}-t_{1}}e^{\sigma\Delta}g\,d\sigma$
for any $g\in L^{2}(\mathcal{D})$, see [22, Proposition 2.1.4]. Applying $(-\Delta)^{-\gamma}$, using $1-\gamma<1$, the fact that $(-\Delta)^{1-\gamma}$ is a closed operator and $\sigma\mapsto(-\Delta)^{1-\gamma}e^{\sigma\Delta}$ is integrable for any $g\in L^{2}(\mathcal{D})$, we obtain, thanks to [21, Proposition A.5],
$(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)g=-\int_{0}^{t_{2}-t_{1}}(-\Delta)^{1-\gamma}e^{\sigma\Delta}g\,d\sigma.$
Again by the estimate for the fractional powers, we have
$\begin{split}\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)g\|_{2}&\leq\int_{0}^{t_{2}-t_{1}}\|(-\Delta)^{1-\gamma}e^{\sigma\Delta}g\|_{2}\,d\sigma\\ &\leq C\int_{0}^{t_{2}-t_{1}}\sigma^{-1+\gamma}\|g\|_{2}\,d\sigma\\ &=\frac{C}{\gamma}(t_{2}-t_{1})^{\gamma}\|g\|_{2}.\end{split}$
Since the previous estimate holds for any $g\in L^{2}(\mathcal{D})$, we have checked
(3.8) $\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}\leq\frac{C}{\gamma}(t_{2}-t_{1})^{\gamma}$
for any $\gamma\in(0,\alpha_{\theta}+\beta_{\theta}).$ Substituting (3.7) and (3.8) into (3.6) gives (3.5).
Thus, to estimate $I_{2}$, using the independence of $(\beta_{k})_{k}$, Itô-isometry and (3.5), we obtain, for any $\gamma\in(0,\alpha_{\theta}+\beta_{\theta})$,
(3.9) $\begin{split}\mathbb{E}\left[\|I_{2}\|_{2}^{2}\right]&\leq\sum_{k}\lambda_{k}^{2}\int_{0}^{t_{1}}\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}\left[e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta}\right](-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,dr\\ &\leq C\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,(t_{2}-t_{1})^{2\gamma}\int_{0}^{t_{1}}(t_{1}-r)^{-2(1-\alpha_{\theta}-\beta_{\theta}+\gamma)}\,dr.\end{split}$
Note that the integral on the right-hand side is finite if $1-\alpha_{\theta}-\beta_{\theta}+\gamma<\frac{1}{2},$ i.e. $\gamma<\alpha_{\theta}+\beta_{\theta}-\frac{1}{2}=\frac{\delta_{\theta}}{4}$. Thus, choosing $\gamma\in(0,\frac{\delta_{\theta}}{4})$, we arrive to
(3.10) $\mathbb{E}\left[\|I_{2}\|_{2}^{2}\right]\leq C\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}(t_{2}-t_{1})^{2\gamma},$
where $C>0$ depends on $T,\alpha_{\theta},\beta_{\theta},\gamma$ but not on $t_{1},t_{2}$. In conclusion, applying Jensen’s inequality and setting $\eta:=\min(\delta_{\theta}/2,2\gamma)\in(0,1]$, we obtain (3.2) by combining the estimates for $I_{1}$, $I_{2}$ in (3.4) and (3.10).
We conclude this section by recalling, and then applying, the notion of stochastic convolution to prove point (iii) of Proposition 2.1.
Given $U,H$, Hilbert spaces, let $U_{0}:=\mathcal{Q}^{1/2}(U)$ and set $L^{0}_{2}=L_{2}(U_{0},H)$ the space of Hilbert-Schmidt operators from $U_{0}$ to $H$, with norm
$\norm{R}_{L^{0}_{2}}=\norm{R\circ\mathcal{Q}^{1/2}}_{L_{2}(U,H)},$
where $L_{2}=L_{2}(U,H)$ is the space of Hilbert-Schmidt operators from $U$ to $H$ with norm
$\norm{R}_{L_{2}}^{2}=\sum_{k}\norm{Re_{k}}_{H}^{2},$
where again $(e_{k})_{k}\subset U$ denotes an orthonormal basis of $U$. Consider $(W_{t})_{t}$ to be a $U$-valued $\mathcal{Q}$-Wiener process and $\mathcal{A}\colon D(\mathcal{A})\subset H\to H$ a linear operator which generates a $C_{0}$-semigroup $(e^{t\mathcal{A}})_{t\geq 0}$ in $H$. Then, we consider the following space of stochastic processes
$\mathcal{N}_{W}^{2}(0,T):=\left\{\Phi\colon[0,T]\times\Omega\to L^{0}_{2}\;\mid\;\Phi\text{ is predictable and }\norm{\Phi}_{T}<\infty\right\},$
where
$\norm{\Phi}_{T}:=\left(\mathbb{E}\left[\int_{0}^{T}\norm{\Phi(t)}_{L^{0}_{2}}^{2}\,dt\right]\right)^{\frac{1}{2}}.$
Lastly, for $\Phi\in\mathcal{N}^{2}_{W}(0,T)$ we recall that the stochastic convolution $W_{\mathcal{A}}^{\Phi}$ is defined as
$W_{\mathcal{A}}^{\Phi}(t)=\int_{0}^{t}e^{(t-r)\mathcal{A}}\Phi(r)\,dW_{r},\quad t\in[0,T],$
and we refer to [11] for more information. We can now move to the proof of the last part of Proposition 2.1.
Proof of Proposition 2.1 (iii).
By Markov’s inequality, for any $\vartheta>0$, we have
(3.11) $\begin{split}\mathbb{P}\left(\sup_{0\leq t\leq T}\norm{Z_{t}^{\varepsilon}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}>\vartheta\right)&\leq\frac{1}{\vartheta^{2}}\mathbb{E}\left[\sup_{0\leq t\leq T}\norm{Z_{t}^{\varepsilon}}^{2}_{H^{-2\alpha_{\theta}}(\mathcal{D})}\right].\end{split}$
To obtain the thesis, we thus need to bound the expectation on the right-hand side. We start to observe that, since $Z_{t}^{\varepsilon}=(-\Delta)^{\alpha_{\theta}}\xi_{t}^{\varepsilon}$, then
$\norm{Z^{\varepsilon}_{t}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}=\norm{(-\Delta)^{\alpha_{\theta}}\xi_{t}^{\varepsilon}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}\cong\norm{\xi_{t}^{\varepsilon}}_{2}.$
Second, recall that
$\xi_{t}^{\varepsilon}=\sqrt{\varepsilon}\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{(t-r)\Delta}(-\Delta)^{\beta_{\theta}}D\,dW_{r}.$
Set $\Phi:=(-\Delta)^{\beta_{\theta}}D$, and observe that $\Phi\in L^{0}_{2}$ since
$\|\Phi\|_{L^{0}_{2}}^{2}=\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}<\infty,$
where $(e_{k})_{k}$ is an orthonormal basis of $U$ made of eigenvectors for the covariance operator $\mathcal{Q}.$ Thus, it is well-defined the $L^{2}(\mathcal{D})$-stochastic convolution $W^{\Phi}_{\Delta}(t):=\int_{0}^{t}e^{(t-r)\Delta}\Phi\,dW_{r}$, for $t\in[0,T]$, and $\xi_{t}^{\varepsilon}=\sqrt{\varepsilon}(-\Delta)^{\frac{1}{2}-\frac{\delta_{\theta}}{4}}W_{\Delta}^{\Phi}(t)$. Choose now $\rho\in(\frac{1}{2}-\frac{\delta_{\theta}}{4},\frac{1}{2})$, and set
$Y_{\rho}(t):=\int_{0}^{t}(t-r)^{-\rho}e^{(t-r)\Delta}\Phi dW_{r},\quad t\in[0,T].$
By the factorisation method, see [11, Section 5.3.1], since $\int_{0}^{T}t^{-2\rho}\|e^{t\Delta}\Phi\|_{L^{0}_{2}}^{2}dt<\infty$, it holds
$W^{\Phi}_{\Delta}(t)=\frac{\sin(\pi\rho)}{\pi}\int_{0}^{t}(t-r)^{\rho-1}e^{(t-r)\Delta}Y_{\rho}(r)\,dr.$
Thus, using the estimate for the fractional powers as in (3.7), we have
$\begin{split}\|\xi_{t}^{\varepsilon}\|_{2}\leq C\sqrt{\varepsilon}\int_{0}^{t}(t-r)^{\rho-1}\|(-\Delta)^{\frac{1}{2}-\frac{\delta_{\theta}}{4}}e^{(t-r)\Delta}\|_{\mathcal{L}(L^{2})}\|Y_{\rho}(r)\|_{2}dr\leq C\sqrt{\varepsilon}\int_{0}^{t}(t-r)^{\rho-\frac{3}{2}+\frac{\delta_{\theta}}{4}}\|Y_{\rho}(r)\|_{2}dr.\end{split}$
Fix now $q>2$ such that $\frac{1}{q}<\rho-(\frac{1}{2}-\frac{\delta_{\theta}}{4})$, and let $q^{\prime}$ be the conjugate exponent of $q$, i.e. $q^{\prime}=\frac{q}{q-1}.$ Then $g(\sigma):=\sigma^{\rho-1-(\frac{1}{2}-\frac{\delta_{\theta}}{4})}\in L^{q^{\prime}}(0,T)$, and by Hölder inequality
$\sup_{0\leq t\leq T}\|\xi^{\varepsilon}_{t}\|_{2}\leq C\sqrt{\varepsilon}\|g\|_{L^{q^{\prime}}(0,T)}\|Y_{\rho}\|_{L^{q}(0,T;L^{2}(\mathcal{D}))}.$
Considering the $q$-th moment, we have
(3.12) $\mathbb{E}\left[\sup_{0\leq t\leq T}\|\xi_{t}^{\varepsilon}\|_{2}^{q}\right]\leq C\varepsilon^{q/2}\mathbb{E}\left[\int_{0}^{T}\|Y_{\rho}(r)\|_{2}^{q}\,dr\right].$
Since $Y_{\rho}$ is a centered Gaussian random variable in $L^{2}(\mathcal{D})$, there exists $c_{q}>0$ such that $\mathbb{E}[\|Y_{\rho}(r)\|_{2}^{q}]\leq c_{q}\left(\mathbb{E}[\|Y_{\rho}(r)\|_{2}^{2}]\right)^{q/2}$. By Itô isometry and the contraction of $e^{t\Delta}$ on $L^{2}$, we have
$\mathbb{E}\left[\|Y_{\rho}(r)\|_{2}^{2}\right]=\int_{0}^{r}(r-\sigma)^{-2\rho}\|e^{(r-\sigma)\Delta}\Phi\|_{L^{0}_{2}}^{2}\,d\sigma\leq\|\Phi\|_{L^{0}_{2}}^{2}\int_{0}^{r}(r-\sigma)^{-2\rho}\,d\sigma=\|\Phi\|_{L^{0}_{2}}^{2}\frac{r^{1-2\rho}}{1-2\rho}.$
Therefore
(3.13) $\mathbb{E}\left[\int_{0}^{T}\|Y_{\rho}(r)\|_{2}^{q}\,dr\right]\leq c_{q}\|\Phi\|_{L^{0}_{2}}^{q}\int_{0}^{T}r^{\frac{(1-2\rho)q}{2}}dr\leq c_{q,\rho,T}\|\Phi\|_{L^{0}_{2}}^{q}.$
Thus, substituting (3.13) inside (3.12), we obtain
$\mathbb{E}\left[\sup_{0\leq t\leq T}\|\xi_{t}^{\varepsilon}\|_{2}^{q}\right]\leq C(\delta_{\theta},T,q)\varepsilon^{q/2}\|\Phi\|_{L^{0}_{2}}^{q}.$
Lastly, since $q>2$, by Jensen inequality we deduce
$\begin{split}\mathbb{E}\left[\sup_{0\leq t\leq T}\|\xi_{t}^{\varepsilon}\|_{2}^{2}\right]\leq\left(\mathbb{E}\left[\sup_{0\leq t\leq T}\|\xi_{t}^{\varepsilon}\|_{2}^{q}\right]\right)^{2/q}\leq\varepsilon C(\delta_{\theta},T)\|\Phi\|_{L^{0}_{2}}^{2}=\varepsilon C(\delta_{\theta},T)\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}.\end{split}$
The thesis then follows by substituting the previous bound into (3.11). ∎
3.2. The remainder temperature equation
Let $\delta_{u}>0$ be a small regularity exponent for the velocity. Given a prescribed divergence-free velocity field
$u\in L^{p}(0,T;H^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))\cap L^{\infty}(0,T;H^{\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})),\quad p\in(1,\infty)$
in this section we consider the remainder $\zeta_{t}^{\varepsilon}:=\theta^{\varepsilon}_{t}-Z^{\varepsilon}_{t}$ between the solution $\theta_{t}^{\varepsilon}$ of the system (2.2), and the solution $Z^{\varepsilon}_{t}$ of the linear problem (2.4) considered in the previous section.
We start by proving an auxiliary interpolation result for the velocity field $u$, which will be used throughout this section.
Corollary 3.1.
Let $\delta_{u}\in(0,1)$ and $p\in(1,\infty)$. If
$u\in L^{p}(0,T;H^{\frac{3}{2}-\delta_{u}}(\mathcal{D}))\cap L^{\infty}(0,T;H^{\frac{1}{2}-\delta_{u}}(\mathcal{D})),$
then for all $0<\lambda\leq 1-\delta_{u}$, we have
$u\in L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T;H^{\frac{1}{2}+\lambda}(\mathcal{D})\right).$
Proof.
Note that if $\lambda=1-\delta_{u}$, then the statement is trivial. Thus, fix $0<\lambda<1-\delta_{u}$, and define $\delta^{\prime}:=1-\lambda.$ Then $\delta_{u}<\delta^{\prime}<1+\delta_{u}$. Since $\frac{3}{2}-\delta^{\prime}\in(\frac{1}{2}-\delta_{u},\frac{3}{2}-\delta_{u})$, there exists $\vartheta\in(0,1)$ such that
$\frac{3}{2}-\delta^{\prime}=\left(1-\vartheta\right)\left(\frac{1}{2}-\delta_{u}\right)+\vartheta\left(\frac{3}{2}-\delta_{u}\right),$
which yields
$\vartheta=1-\delta^{\prime}+\delta_{u}.$
Using the interpolation inequality
$\left\|u_{t}\right\|_{H^{\frac{3}{2}-\delta^{\prime}}(\mathcal{D})}\leq C\left\|u_{t}\right\|_{H^{\frac{1}{2}-\delta_{u}}(\mathcal{D})}^{\delta^{\prime}-\delta_{u}}\|u_{t}\|_{H^{\frac{3}{2}-\delta_{u}}(\mathcal{D})}^{1-\delta^{\prime}+\delta_{u}},$
for a constant $C>0$. Therefore, for any $q\geq 1$, we have
$\int_{0}^{T}\left\|u_{t}\right\|_{H^{\frac{3}{2}-\delta^{\prime}}(\mathcal{D})}^{q}dt$ $\leq C\int_{0}^{T}\left\|u_{t}\right\|_{H^{\frac{1}{2}-\delta_{u}}(\mathcal{D})}^{\left(\delta^{\prime}-\delta_{u}\right)q}\left\|u_{t}\right\|_{H^{\frac{3}{2}-\delta_{u}}(\mathcal{D})}^{\left(1-\delta^{\prime}+\delta_{u}\right)q}dt$
$\leq C\left\|u\right\|_{L^{\infty}\left(0,T;H^{\frac{1}{2}-\delta_{u}}(\mathcal{D})\right)}^{\left(\delta^{\prime}-\delta_{u}\right)q}\left\|u\right\|_{L^{\left(1-\delta^{\prime}+\delta_{u}\right)q}\left(0,T;H^{\frac{3}{2}-\delta_{u}}(\mathcal{D})\right)}^{\left(1-\delta^{\prime}+\delta_{u}\right)q}.$
This is finite if $(1-\delta^{\prime}+\delta_{u})q=p$, which gives
$q=\frac{p}{1-\delta^{\prime}+\delta_{u}}=\frac{p}{\lambda+\delta_{u}}.$
Further, observe that $\frac{3}{2}-\delta^{\prime}=\lambda+\frac{1}{2}$ by definition of $\lambda$. Thus, we have proved
$\int_{0}^{T}\norm{u_{t}}_{H^{\lambda+\frac{1}{2}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}<\infty,$
which gives the thesis. ∎
To make sense of the mild solution, we next establish a space-time integrability result for the product $u_{t}\cdot Z_{t}^{\varepsilon}$. Recall that, for any $\delta_{\theta}>0$, $\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2}$ and $\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4}$, we have introduced the stochastic processes
$Z_{t}^{\varepsilon}=(-\Delta)^{\alpha_{\theta}}\xi_{t}^{\varepsilon},\quad\xi_{t}^{\varepsilon}=\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}}e^{(t-r)\Delta}DdW_{r}.$
Further, if
$\sum_{k}\lambda_{k}^{2}\norm{(-\Delta)^{\beta_{\theta}}De_{k}}_{2}^{2}<\infty,$
then by Proposition 2.1 we know that the trajectories of $\xi^{\varepsilon}$ are in $C(0,T;L^{2}(\mathcal{D}))$.
Lemma 3.2.
Let $0<\lambda<1$ and $\gamma>\frac{1}{4}$. Set
$\overline{\delta}_{\theta}(\lambda,\gamma):=\min\{\lambda,2\gamma-\frac{1}{2}\}.$
For any $0<\delta_{\theta}\leq\overline{\delta}_{\theta}$, any velocity field
$u\in L^{p}(0,T;H^{\frac{3}{2}-\delta_{u}}(\mathcal{D}))\cap L^{\infty}(0,T;H^{\frac{1}{2}-\delta_{u}}(\mathcal{D})),\qquad p>1,\,\delta_{u}\in(0,1-\lambda],$
we have
$t\mapsto(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}(-\Delta)^{\alpha_{\theta}}\xi_{t}^{\varepsilon}\right)\in L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T;L^{\frac{6}{5}}(\mathcal{D})\right),\qquad\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2}.$
Moreover,
(3.14) $\norm{(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}(u_{t}(-\Delta)^{\alpha_{\theta}}\xi_{t}^{\varepsilon})}_{L^{\frac{p}{\lambda+\delta_{u}}}(0,T;L^{\frac{6}{5}}(\mathcal{D}))}\leq C\norm{\xi^{\varepsilon}}_{C(0,T;L^{2}(\mathcal{D}))}\,\norm{u}_{L^{\frac{p}{\lambda+\delta_{u}}}(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))},$
where $C=C(\lambda,\gamma,\delta_{\theta})>0$.
Proof.
To simplify the notation, we set $\varepsilon=1$ and drop the dependence on $\varepsilon$ of $\xi^{\varepsilon}$. We start by recalling that from [6, Theorem 7.4], given $s\geq 0$, the pointwise multiplication
$W^{s_{1},p_{1}}\left(\mathcal{D}\right)\times W^{s_{2},p_{2}}\left(\mathcal{D}\right)$ $\hookrightarrow W^{s,p}\left(\mathcal{D}\right)$
$(f,g)$ $\mapsto fg$
is bilinear and continuous if
  1. (i)
    $s_{1}\geq s$ and $s_{2}\geq s$;
  2. (ii)
    $s_{1}\geq s+3\left(\frac{1}{p_{1}}-\frac{1}{p}\right)$ and $s_{2}\geq s+3\left(\frac{1}{p_{2}}-\frac{1}{p}\right)$;
  3. (iii)
    $s_{1}+s_{2}>s+3\left(\frac{1}{p_{1}}+\frac{1}{p_{2}}-\frac{1}{p}\right).$
Step 1. As a first step of the proof, we show that
$\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}(-\Delta)^{\alpha_{\theta}}\xi_{t}\right),\varphi\right\rangle\leq K_{t}\left\|\varphi\right\|_{L^{6}\left(\mathcal{D}\right)},$
where the integrability in time of $t\mapsto K_{t}$ will be investigated in Step $2$. Here, $6$ is the conjugate exponent of $\frac{6}{5}$, $\varphi\in L^{6}(\mathcal{D})$ is a test function, and $\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2}$. Further, we denote by $\langle\cdot,\cdot\rangle$ the duality pairing. Since $\Delta$ is self-adjoint, by Hölder inequality and integration by parts, we obtain
$\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}(-\Delta)^{\alpha_{\theta}}\xi_{t}\right),\varphi\right\rangle$ $=\left\langle\mathrm{div}(u_{t}(-\Delta)^{\alpha_{\theta}}\xi_{t}),(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle$
$=-\langle u_{t}(-\Delta)^{\alpha_{\theta}}\xi_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle$
$=-\left\langle(-\Delta)^{\alpha_{\theta}}\xi_{t},u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle$
$=-\left\langle\xi_{t},(-\Delta)^{\alpha_{\theta}}\left(u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right)\right\rangle$
$\leq\left\|\xi_{t}\right\|_{L^{2}\left(\mathcal{D}\right)}\left\|(-\Delta)^{\alpha_{\theta}}\left(u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right)\right\|_{L^{2}\left(\mathcal{D}\right)}$
$=\left\|\xi_{t}\right\|_{L^{2}\left(\mathcal{D}\right)}\left\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{W^{2\alpha_{\theta},2}\left(\mathcal{D}\right)}.$
Next, we show that $u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\in W^{2\alpha_{\theta},2}\left(\mathcal{D}\right)$. Since $2\alpha_{\theta}=\frac{1}{2}+\delta_{\theta}$, this is equivalent to
(3.15) $u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\in W^{\frac{1}{2}+\delta_{\theta},2}\left(\mathcal{D}\right).$
By the norm equivalence between $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})$ and $H^{\frac{1}{2}+\lambda}(\mathcal{D})$, and by Corollary 3.1, we have $u_{t}\in W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)$ a.e., and since $\varphi\in L^{6}\left(\mathcal{D}\right)$, we deduce $\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\in W^{2\gamma,6}\left(\mathcal{D}\right)$. Hence, (3.15) follows from the pointwise multiplication embedding
$W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)\times W^{2\gamma,6}\left(\mathcal{D}\right)$ $\hookrightarrow W^{\frac{1}{2}+\delta_{\theta},2}\left(\mathcal{D}\right)$
$(u_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi)$ $\mapsto u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi,$
provided that the following conditions hold
  1. (i)
    $\frac{1}{2}+\lambda\geq\frac{1}{2}+\delta_{\theta}$ and $2\gamma\geq\frac{1}{2}+\delta_{\theta}$;
  2. (ii)
    $\frac{1}{2}+\lambda\geq\frac{1}{2}+\delta_{\theta}+3\left(\frac{1}{2}-\frac{1}{2}\right)$ and $2\gamma\geq\frac{1}{2}+\delta_{\theta}+3\left(\frac{1}{6}-\frac{1}{2}\right)$;
  3. (iii)
    $\frac{1}{2}+\lambda+2\gamma>\frac{1}{2}+\delta_{\theta}+3\left(\frac{1}{2}+\frac{1}{6}-\frac{1}{2}\right)$.
The previous conditions are equivalent to
  1. (a)
    $\delta_{\theta}\leq\lambda$ and $\delta_{\theta}\leq 2\gamma-\frac{1}{2}$;
  2. (b)
    $\delta_{\theta}\leq\frac{1}{2}+2\gamma$;
  3. (c)
    $\delta_{\theta}<\lambda+2\gamma-\frac{1}{2}.$
Therefore, we can choose
(3.16) $\overline{\delta}_{\theta}(\lambda,\gamma):=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\}>0,$
and note that $\overline{\delta}_{\theta}$ is positive since $\gamma>\frac{1}{4}$ and $\lambda\in(0,1).$
Step 2. Fix now $0<\delta_{\theta}\leq\overline{\delta}_{\theta}(\lambda,\gamma)$. The computations of the previous step yield
$\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}(-\Delta)^{\alpha_{\theta}}\xi_{t}\right)\right\|_{L^{\frac{6}{5}}\left(\mathcal{D}\right)}$ $=\sup_{\|\varphi\|_{L^{6}(\mathcal{D})}\leq 1}\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}(-\Delta)^{\alpha_{\theta}}\xi_{t}\right),\varphi\right\rangle$
$\leq C\left\|\xi_{t}\right\|_{L^{2}\left(\mathcal{D}\right)}\sup_{\|\varphi\|_{6}\leq 1}\left\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{W^{2\alpha_{\theta},2}\left(\mathcal{D}\right)}$
$\leq C\left\|\xi_{t}\right\|_{L^{2}\left(\mathcal{D}\right)}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)}\sup_{\|\varphi\|_{6}\leq 1}\left\|\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{W^{2\gamma,6}\left(\mathcal{D}\right)}$
$\leq C\left\|\xi_{t}\right\|_{L^{2}\left(\mathcal{D}\right)}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)}\sup_{\|\varphi\|_{6}\leq 1}\left\|\varphi\right\|_{L^{6}\left(\mathcal{D}\right)}$
$\leq C\left\|\xi_{t}\right\|_{L^{2}\left(\mathcal{D}\right)}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)},$
where $C=C(\lambda,\gamma,\delta_{\theta})$ is a constant from the multiplication results. Thus, up to positive constants, we obtain from Hölder’s inequality that
$\int_{0}^{T}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}(-\Delta)^{\alpha_{\theta}}\xi_{t}\right)\right\|_{L^{\frac{6}{5}}\left(\mathcal{D}\right)}^{\frac{p}{\lambda+\delta_{u}}}dt$ $\leq\int_{0}^{T}\left(\left\|\xi_{t}\right\|_{L^{2}\left(\mathcal{D}\right)}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)}\right)^{\frac{p}{\lambda+\delta_{u}}}dt$
$=\int_{0}^{T}\left\|\xi_{t}\right\|_{L^{2}\left(\mathcal{D}\right)}^{\frac{p}{\lambda+\delta_{u}}}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)}^{\frac{p}{\lambda+\delta_{u}}}dt$
$\leq\sup_{0\leq t\leq T}\left\|\xi_{t}\right\|_{L^{2}\left(\mathcal{D}\right)}^{\frac{p}{\lambda+\delta_{u}}}\int_{0}^{T}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)}^{\frac{p}{\lambda+\delta_{u}}}dt.$
Note that all the terms on the right-hand side are finite thanks to Proposition 2.1, Corollary 3.1, and the norm equivalence between $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})$ and $H^{\frac{1}{2}+\lambda}(\mathcal{D})$; the last corollary can be applied since by hypothesis $\delta_{u}\leq 1-\lambda$. ∎
Recall that the mild solution of the remainder’s problem (2.5) has the form
$\zeta_{t}^{\varepsilon}=e^{t\Delta}\theta_{0}-\int_{0}^{t}e^{\left(t-r\right)\Delta}\left(u_{r}^{\varepsilon}\cdot\nabla\zeta_{r}^{\varepsilon}\right)dr-\int_{0}^{t}e^{\left(t-r\right)\Delta}\left(u_{r}^{\varepsilon}\cdot\nabla Z_{r}^{\varepsilon}\right)dr.$
By the previous auxiliary results, we can deduce the regularity of the third term on the right-hand side of the previous identity.
Corollary 3.3.
Let $s\in[0,\frac{1}{2})$, $p>1$ and $\gamma>\frac{1}{4}$ be given. Assume that
(3.17) $a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}<1.$
Let $\lambda\in(0,1)$ such that
(3.18) $\lambda<p(1-a(s,\gamma)).$
Fix $0<\delta_{\theta}\leq\overline{\delta}_{\theta}=\min\{\lambda,2\gamma-\frac{1}{2}\},$ and consider $\alpha_{\theta}:=\frac{1}{4}+\frac{\delta_{\theta}}{2}.$ Define
$\overline{\delta}_{u}(p,s,\gamma,\lambda):=\min\{p(1-a(s,\gamma))-\lambda,1-\lambda\}>0.$
Then, for every $\delta_{u}\in(0,\overline{\delta}_{u})$, and every velocity field
$u\in L^{p}(0,T;H^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))\cap L^{\infty}(0,T;H^{\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})),$
the function
$\chi_{t}:=\int_{0}^{t}e^{(t-r)\Delta}\bigl(u_{r}\cdot\nabla Z_{r}^{\varepsilon}\bigr)\,dr$
belongs to $C(0,T;W^{s,\frac{6}{5}}(\mathcal{D}))$ and satisfies
(3.19) $\|\chi^{\varepsilon}\|_{C(0,T;W^{s,\frac{6}{5}})}\leq C\,\|\xi^{\varepsilon}\|_{C(0,T;L^{2})}\,\|u\|_{L^{\frac{p}{\lambda+\delta_{u}}}(0,T;W^{\frac{1}{2}+\lambda,2})},$
where $C=C(s,p,\gamma,\lambda,\delta_{\theta},T)>0$. Lastly, given $s\in[0,\frac{1}{2}),p>1$ as before, there always exists $\gamma=\gamma(s)>\frac{1}{4}$ and $\lambda=\lambda(s,p)\in(0,1)$ such that (3.17) and (3.18) hold.
Proof.
To simplify the notation, we set $\varepsilon=1$ and drop the dependence of $\varepsilon$ of $\xi^{\varepsilon}$.
Step 1. Fix $\delta_{\theta}>0$ such that $0<\delta_{\theta}<\overline{\delta}_{\theta}=\min\{\lambda,2\gamma-\frac{1}{2}\}$. Define the function
$f_{\lambda}(\delta_{u})=(\gamma+\frac{s}{2}+\frac{1}{2})\frac{1}{1-\frac{\lambda}{p}-\frac{\delta_{u}}{p}}=\frac{a(s,\gamma)}{1-\frac{\lambda+\delta_{u}}{p}},\quad\lambda\geq 0.$
If $\delta_{u}<\overline{\delta}_{u}$, then $\lambda+\delta_{u}<p(1-a(s,\gamma))$, which implies
(3.20) $f_{\lambda}(\delta_{u})=\bigl(\gamma+\tfrac{s}{2}+\tfrac{1}{2}\bigr)\,\frac{1}{1-\frac{\lambda}{p}-\frac{\delta_{u}}{p}}<1.$
Step 2. Thanks to the commuting properties of the fractional Laplacian with the heat semigroup and the divergence free condition of $u$, it holds
$\chi_{t}$ $=(-\Delta)^{-\frac{s}{2}}\rho_{t},$
$\rho_{t}$ $=\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}(-\Delta)^{\alpha_{\theta}}\xi_{r}\right)dr.$
By the estimate
$\norm{(-\Delta)^{\varsigma}e^{(t-r)\Delta}y}_{L^{\frac{6}{5}}(\mathcal{D})}\leq\frac{c}{(t-r)^{\varsigma}}\norm{y}_{L^{\frac{6}{5}}(\mathcal{D})},$
where $\varsigma:=\gamma+\frac{s}{2}+\frac{1}{2}$ and $c=c(\varsigma)$ is a positive constant, we get
(3.21) $\begin{split}\left\|\rho_{t}\right\|_{L^{\frac{6}{5}}(\mathcal{D})}&\leq\int_{0}^{t}\frac{c}{\left(t-r\right)^{\gamma+\frac{s}{2}+\frac{1}{2}}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}(-\Delta)^{\alpha_{\theta}}\xi_{r}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}dr\\ &\leq\left(\int_{0}^{t}\left(\frac{c}{\left(t-r\right)^{\gamma+\frac{s}{2}+\frac{1}{2}}}\right)^{\frac{p}{p-\lambda-\delta_{u}}}dr\right)^{\frac{p-\lambda-\delta_{u}}{p}}\left(\int_{0}^{t}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}(-\Delta)^{\alpha_{\theta}}\xi_{r}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dr\right)^{\frac{\lambda+\delta_{u}}{p}}.\end{split}$
The second term is finite thanks to Lemma 3.2 (that we can apply since $\lambda\leq 1-\delta_{u}$ and $\delta_{\theta}\in(0,\overline{\delta}_{\theta}]$), while the first term is finite thanks to (3.20). Taking the supremum over $t\in[0,T]$ in (3.21), using Lemma 3.2 to bound the second term on the right-hand side of (3.21), and the boundedness of $(-\Delta)^{-s/2}\colon L^{6/5}\to W^{s,6/5}$, we obtain (3.19).
Step 3. We now prove that $t\mapsto\rho_{t}$ is continuous as a map $(0,T)\to L^{\frac{6}{5}}(\mathcal{D})$. For notational convenience, write
$\rho_{t}=\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(t-\tau)\Delta}g_{\tau}\,d\tau,\qquad g_{\tau}:=(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\bigl(u_{\tau}(-\Delta)^{\alpha_{\theta}}\xi_{\tau}\bigr),$
so that $g\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T;L^{\frac{6}{5}}(\mathcal{D}))$ by Lemma 3.2. Fix $r\in(0,T]$ and let $t\in[r,T]$. Then
$\rho_{t}-\rho_{r}$ $=\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(t-\tau)\Delta}g_{\tau}\,d\tau-\int_{0}^{r}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(r-\tau)\Delta}g_{\tau}\,d\tau$
$=\int_{r}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(t-\tau)\Delta}g_{\tau}\,d\tau+\int_{0}^{r}\Bigl[(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(t-\tau)\Delta}-(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(r-\tau)\Delta}\Bigr]g_{\tau}\,d\tau$
$=:I_{1}(t,r)+I_{2}(t,r).$
Using the estimate for the fractional powers as in (3.7) and Hölder’s inequality with $\frac{p}{\lambda+\delta_{u}}>1, q=q$$^{\prime}=\frac{q}{q-1}$, we obtain
$\|I_{1}(t,r)\|_{L^{\frac{6}{5}}}\leq C\Bigl(\int_{r}^{t}(t-\tau)^{-\bigl(\gamma+\frac{s}{2}+\frac{1}{2}\bigr)q^{\prime}}\,d\tau\Bigr)^{1/q^{\prime}}\Bigl(\int_{r}^{t}\|g_{\tau}\|_{L^{\frac{6}{5}}}^{q}\,d\tau\Bigr)^{1/q},$
for $C=C(\gamma,s,p,\lambda,\delta_{u})>0$. The first factor $I_{1}$ is bounded uniformly in $t\in[r,T]$ thanks to condition (3.20), while the second factor tends to 0 as $t\downarrow r$ by the absolute continuity of the integral, since $g\in L^{q}(0,T;L^{\frac{6}{5}}(\mathcal{D}))$. Hence $\|I_{1}(t,r)\|_{L^{\frac{6}{5}}}\to 0$ as $t\downarrow r$.
For the second term $I_{2}$, for every fixed $\tau\in(0,r)$ the map
$\theta\mapsto(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\theta\Delta}g_{\tau},\qquad\theta>0,$
is continuous in $L^{\frac{6}{5}}(\mathcal{D})$ by strong continuity of the heat semigroup; therefore the integrand in $I_{2}(t,r)$ converges to 0 in $L^{\frac{6}{5}}$ as $t\downarrow r$ for a.e. $\tau\in(0,r)$. Moreover, using again the estimate for the fractional power (3.7), for $t$ close to $r$ we have
$\|(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(t-\tau)\Delta}g_{\tau}-(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(r-\tau)\Delta}g_{\tau}\|_{L^{\frac{6}{5}}}\leq C(r-\tau)^{-\gamma-\frac{s}{2}-\frac{1}{2}}\|g_{\tau}\|_{L^{\frac{6}{5}}},$
and the right-hand side is integrable on $(0,r)$ by the same computation as in (3.21). Hence, by the dominated convergence theorem, $\|I_{2}(t,r)\|_{L^{\frac{6}{5}}}\to 0$ as $t\downarrow r$. We deduce that $\|\rho_{t}-\rho_{r}\|_{L^{\frac{6}{5}}(\mathcal{D})}\to 0$ for every $r\in(0,T]$, that is $\rho\in C(0,T;L^{\frac{6}{5}}(\mathcal{D}))$. Consequently, $\chi_{t}=(-\Delta)^{-\frac{s}{2}}\rho_{t}$ belongs to $C(0,T;W^{s,\frac{6}{5}}(\mathcal{D})\bigr)$.
Step 4. Existence of $\gamma,\lambda$, given $s,p$. Assume that, as before, $s\in[0,\frac{1}{2})$ and $p>1$ are given. We now check that we can always find $\gamma=\gamma(s)>\frac{1}{4}$ and $\lambda=\lambda(s,p)$ such that (3.17) and (3.18) are satisfied, and thus this corollary is non-trivial. Indeed, with the notation of Step $1$, from $a(\frac{1}{4})=\frac{3}{4}+\frac{s}{2}<1$, it follows that there exists $\gamma=\gamma(s)>\frac{1}{4}$ sufficiently close to $\frac{1}{4}$ such that $a(s,\gamma)<1$, that is (3.17). Hence, also p(1a)>0p(1-a)>0 holds, and we may choose $\lambda=\lambda(s,p)\in(0,1)$ so small that $\lambda<p(1-a),$ that is (3.18). ∎
Remark 3.4.
From the proof of Corollary 3.3, it follows that if $s\in[0,\frac{1}{2}),p>1$, $\gamma>\frac{1}{4}$ and $\lambda\in(0,1)$ satisfy (3.17) and (3.18), then
$\bigl(\gamma+\tfrac{s}{2}+\tfrac{1}{2}\bigr)\,\frac{1}{1-\frac{\lambda}{p}-\frac{\delta_{u}}{p}}<1$
for any $\delta_{u}<\overline{\delta}_{u}=\min\{p(1-a(s,\gamma))-\lambda,1-\lambda\}$.
Lastly, we show the proof of Theorem 2.2.
Proof of Theorem 2.2.
For simplicity we set $\varepsilon=1$ and drop the superscript, writing $\zeta$ and $Z$ instead of $\zeta^{\varepsilon}$ and $Z^{\varepsilon}$. We also use the notation of Corollary 3.3 and Proposition 2.1. Let $s\in[0,\frac{1}{2})$ and $p>2$ be given.
Step $1$. Choices of $\gamma,q,q^{\prime},\lambda,\overline{\delta}_{\theta},\overline{\delta}_{u}.$
Choice of $\gamma=\gamma(s,p)>\frac{1}{4}.$ We choose $\gamma=\frac{1}{4}+\varepsilon_{\gamma}$, with $\varepsilon_{\gamma}>0$ sufficiently small such that
(3.22) $\left\{\begin{aligned} a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}&<1,\\ \lambda_{\max}(s,p,\gamma):=p(1-a(s,\gamma))&>\frac{1}{2}-s.\end{aligned}\right$.
The condition (3.22)1 is satisfied, up to choosing $\varepsilon_{\gamma}$ sufficiently small, since
$\lim\limits_{\gamma\downarrow\frac{1}{4}}a(s,\gamma)=\frac{3}{4}+\frac{s}{2}<1,$
where in the last inequality we use $s<\frac{1}{2}$. For the condition (3.22)2, note that if $\gamma\in(\frac{1}{4},\frac{1-s}{2})$, then $0<\frac{1}{2}-\gamma-\frac{s}{2}=1-a(s,\gamma)$. Further, since $p>2$ by hypothesis, we have
$\lim\limits_{\gamma\downarrow\frac{1}{4}}\lambda_{\max}(s,p,\gamma)=\frac{p}{2}(\frac{1}{2}-s)>\frac{1}{2}-s.$
Choice of $q=q(s)$ and $q^{\prime}=q^{\prime}(s)$. Let $q=q(s):=\frac{6}{5-2s}.$ With this choice, the Sobolev embedding
(3.23) $W^{s,\frac{6}{5}}(\mathcal{D})\hookrightarrow L^{q(s)}(\mathcal{D})$
holds, a condition that we will later need in the proof. Further, denoting by $q^{\prime}$ the conjugate exponent of $q$, it holds
(3.24) $1-\frac{3}{q^{\prime}}<\lambda_{\max}(s,\gamma,p).$
Indeed
$1-\frac{3}{q^{\prime}(s)}=-2+\frac{3}{q(s)}=\frac{1}{2}-s,$
and (3.24) then follows from (3.22)2.
Choice of $\lambda=\lambda(s).$ Define
$\lambda_{\min}(s):=\frac{1}{2}-s=1-\frac{3}{q^{\prime}(s)}.$
and note that $\lambda_{\min}\in(0,1)$ since $s\in[0,\frac{1}{2})$. Further, $\lambda_{\min}<\lambda_{\max}$. Thus, we can choose $\lambda$ sufficiently close to $\lambda_{\min}$ such that
$\lambda=\lambda(s)\in(\lambda_{\min},\lambda_{\max})\cap(0,1).$
With this choice of $\lambda$, we claim that the pointwise multiplication embedding
(3.25) $W^{\frac{1}{2}+\lambda(s),2}(\mathcal{D})\times W^{2\gamma(s),6}(\mathcal{D})\hookrightarrow L^{q^{\prime}(s)}(\mathcal{D})$
holds, another condition that we will need later in the proof. Indeed, thanks to [6, Theorem 7.4], it is sufficient to check that
  1. (i)
    $\frac{1}{2}+\lambda\geq 0$, $2\gamma\geq 0$;
  2. (ii)
    $\frac{1}{2}+\lambda-\frac{3}{2}\geq-\frac{3}{q^{\prime}}$, $2\gamma-\frac{1}{2}\geq-\frac{3}{q^{\prime}}$;
  3. (iii)
    $\frac{1}{2}+\lambda+2\gamma-3(\frac{1}{2}+\frac{1}{6})>-\frac{3}{q^{\prime}}.$
The previous conditions are equivalent to
  1. (a)
    $\lambda\geq-\frac{1}{2},$ $\gamma\geq 0$;
  2. (b)
    $\lambda\geq 1-\frac{3}{q^{\prime}},$ $2\gamma\geq\frac{1}{2}-\frac{3}{q^{\prime}};$
  3. (c)
    $\lambda+2\gamma\geq\frac{3}{2}-\frac{3}{q^{\prime}}$.
The conditions in (a) and (b) follow from the choice of $\lambda,\gamma$, in particular thanks to $\lambda>\lambda_{\min}>0$ and $\gamma>\frac{1}{4}.$ Lastly, since $\lambda+2\gamma>\lambda+\frac{1}{2}$, then (c) is satisfied if $\lambda+\frac{1}{2}>\frac{3}{2}-\frac{3}{q^{\prime}}$, which is true thanks to the choice $\lambda>\lambda_{\min}.$
Choice of $\overline{\delta}_{\theta}$ and $\overline{\delta}_{u}$. With $\gamma$ and $\lambda$ fixed as above, we define the thresholds
$\overline{\delta}_{\theta}(s,p):=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\},\quad\overline{\delta}_{u}(s,p):=\min\left\{\lambda_{\max}(s,p,\gamma)-\lambda,1-\lambda\right\},$
and note that $\overline{\delta}_{\theta}$ and $\overline{\delta}_{u}$ are the same as in Corollary 3.3. Further, both are positive by the choices made above. We assume $0<\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $0<\delta_{u}<\overline{\delta}_{u}$. Since $\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $\delta_{u}<\overline{\delta}_{u}$, the assumptions of Corollary 3.3 are satisfied. Furthermore, from Remark 3.4, it follows that
(3.26) $(\gamma+\frac{s}{2}+\frac{1}{2})\frac{1}{1-\frac{\lambda}{p}-\frac{\delta_{u}}{p}}<1.$
Step 2. Fixed point argument. Observe that, using the notation introduced in Corollary 3.3, we can write $\zeta_{t}$ in the form
$\zeta_{t}=e^{t\Delta}\theta_{0}-(-\Delta)^{-\frac{s}{2}}\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)dr-\chi_{t},$
Let $X:=C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)$, where $T_{1}\leq T$ will be chosen later, and for $\zeta\in X$ consider $\Psi\left(\zeta\right):=\tilde{\zeta}$, where $\tilde{\zeta}$ is defined as follows
(3.27) $\tilde{\zeta}_{t}:=e^{t\Delta}\theta_{0}-(-\Delta)^{-\frac{s}{2}}\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)dr-\chi_{t}.$
We are going to apply a fixed point argument to obtain the thesis. First, since $\chi_{t}\in X$ thanks to Corollary 3.3, to show that $\Psi(\zeta)\in X$ it is sufficient to prove that the linear application
$\Lambda\left(\zeta\right)_{t}:=\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)dr$
maps $X$ to $C(0,T_{1};L^{\frac{6}{5}}(\mathcal{D}))$. Second, we will show that $\Psi\colon X\to X$ is a contraction. We divide the remaining part of the proof into three steps.
Step 3. We want to show that, for any fixed $0<\delta_{u}<\overline{\delta}_{u}(s,p)$ and any $u$ as in the statement of the theorem, it holds
$(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T_{1};L^{\frac{6}{5}}(\mathcal{D})).$
As in the proof of Lemma 3.2, given $\varphi\in L^{6}(\mathcal{D})$ we have
$\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle$ $=\langle\mathrm{div}(u_{t}\zeta_{t}),(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle$
$=-\langle u_{t}\zeta_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle$
$=-\langle\zeta_{t},u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle.$
We now consider the conjugate exponents $q=q(s),q^{\prime}=q^{\prime}(s)\in(1,+\infty)$ chosen in Step 1. We know that $u\in L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)$ thanks to the norm equivalence between $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})$ and $H^{\frac{1}{2}+\lambda}(\mathcal{D})$, and Corollary 3.1 (note that its hypotheses are satisfied since $\delta_{u}\in(0,1)$, $\lambda\in(0,1-\delta_{u})$). Since the pointwise multiplication embedding
(3.28) $\begin{split}W^{\frac{1}{2}+\lambda(s),2}(\mathcal{D})\times W^{2\gamma(s),6}(\mathcal{D})&\hookrightarrow L^{q^{\prime}(s)}(\mathcal{D})\\ (u_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi)&\mapsto u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\end{split}$
holds thanks to the condition verified after (3.25), we get that there exists $C=C(\lambda(s),\gamma(s))=C(s)>0$ such that
$\left\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{L^{q^{\prime}}(\mathcal{D})}\leq C\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\left\|\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{W^{2\gamma,6}(\mathcal{D})},$
and thus
$\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}$ $=\sup_{\|\varphi\|_{6}\leq 1}\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle$
$\leq C\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\sup_{\|\varphi\|_{6}\leq 1}\left\|\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{W^{2\gamma,6}(\mathcal{D})}$
$\leq C\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\sup_{\|\varphi\|_{6}\leq 1}\left\|\varphi\right\|_{L^{6}(\mathcal{D})}$
$=C\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}.$
Further, since $\zeta\in X=C(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D}))\hookrightarrow C(0,T_{1};L^{q}(\mathcal{D}))$ thanks to (3.23), we can write
(3.29) $\int_{0}^{T_{1}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dt\leq C\left(\sup_{t\in[0,T_{1}]}\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\right)^{\frac{p}{\lambda+\delta_{u}}}\int_{0}^{T_{1}}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dt<\infty.$
The thesis of this step follows at this point since $u\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))$.
Step 4. Repeating the same arguments in Corollary 3.3, we obtain $\Lambda\left(\zeta\right)_{t}\in C\left(0,T;L^{\frac{6}{5}}(\mathcal{D})\right)$, and thus $\Psi(\zeta)=\tilde{\zeta}\in X$.
Step 5. We conclude the proof by showing that $\Psi$ is a contraction from $X$ to itself, up to choosing $T_{1}$ small enough. Consider $\zeta_{1},\zeta_{2}\in X$ and their difference $\zeta:=\zeta_{1}-\zeta_{2}$, then, from
$\Psi(\zeta_{i})=e^{t\Delta}\theta_{0}-(-\Delta)^{-\frac{s}{2}}\Lambda(\zeta_{i})_{t}-\chi_{t},$
we deduce
$\Psi(\zeta_{1})-\Psi(\zeta_{2})=-(-\Delta)^{-\frac{s}{2}}\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)dr.$
Considering the norm, we obtain
(3.30) $\begin{split}\norm{\Psi(\zeta_{1})-\Psi(\zeta_{2})}_{X}&\leq\sup_{t\in[0,T_{1}]}\norm{\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)dr}_{L^{\frac{6}{5}}(\mathcal{D})}\\ &\leq\int_{0}^{T_{1}}\norm{(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(T_{1}-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)}_{L^{\frac{6}{5}}(\mathcal{D})}dr.\end{split}$
Using the same argument applied in Step 2 of the proof of Corollary 3.3, we get
(3.31) $\begin{split}\norm{\Psi(\zeta_{1})-\Psi(\zeta_{2})}_{X}&\leq\int_{0}^{T_{1}}\frac{c}{\left(T_{1}-r\right)^{\gamma+\frac{s}{2}+\frac{1}{2}}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}dr\\ &\leq\left(\int_{0}^{T_{1}}\left(\frac{c}{\left(T_{1}-r\right)^{\gamma+\frac{s}{2}+\frac{1}{2}}}\right)^{\frac{p}{p-\lambda-\delta_{u}}}dr\right)^{\frac{p-\lambda-\delta_{u}}{p}}\\ &\quad\cdot\left(\int_{0}^{T_{1}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\,\mathrm{div}(u_{r}\zeta_{r})\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dr\right)^{\frac{\lambda+\delta_{u}}{p}}.\end{split}$
where $c=c(\gamma(s),s)=c(s)$. The first term is finite thanks to condition (3.26). The second term can be bounded as at the end of Step 3 in (3.29), to obtain
$\begin{split}\left(\int_{0}^{T_{1}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dr\right)^{\frac{\lambda+\delta_{u}}{p}}\leq C\sup_{t\in[0,T_{1}]}\norm{\zeta}_{L^{q}(\mathcal{D})}\norm{u}_{L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)}.\end{split}$
Here, $C=C(\lambda(s),\gamma(s))=C(s)$ is independent of the time variable. In conclusion, we have shown
(3.32) $\norm{\Psi(\zeta_{1})-\Psi(\zeta_{2})}_{C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)}\leq K\,\norm{u}_{L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)}\,\norm{\zeta}_{C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)},$
where we observe that $K$ is independent of the time variable. Thus, since $u\in L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)$, then $\Psi$ is a contraction from $X$ to itself if we choose $T_{1}>0$ sufficiently small. This gives the existence of a solution $\zeta\in C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)$. If $T_{1}<T$, we can iteratively repeat the fixed point argument to get the solution of the thesis $\zeta\in C\left(0,T;W^{s,\frac{6}{5}}(\mathcal{D})\right)$ over the entire time interval $[0,T]$.
Remark 3.5.
Note that Corollary 2.3 follows by applying to the definition of $\tilde{\zeta}_{t}$ in (3.27) the same arguments that were used to prove (3.19) and (3.29).
4. Velocity problem
Let $\delta_{u}>0$ be a small regularity exponent for the velocity. In this section, we consider the velocity problem (2.3), assuming that the temperature is a given forcing term $f^{\theta}\in L^{p}(0,T;H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))$. By exploiting the weak-Stokes operator $A_{w}\colon\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\to\prescript{}{0}{H^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})}$ introduced in (2.1) as
$\langle A_{w}u,v\rangle=\langle\nabla u,\nabla v\rangle_{{H^{\frac{1}{2}-\delta_{u}}},{H^{-\frac{1}{2}+\delta_{u}}}},\quad(u,v)\in\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\times\prescript{}{0}{H^{\frac{1}{2}+\delta_{u}}_{\sigma}(\mathcal{D})},$
then the velocity problem can be reformulated as a Cauchy problem in $\prescript{}{0}{H^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})}$ as follows
[…]
where $P$ denotes the Helmholtz projection, and $u_{0}\in V_{p}^{\delta_{u}}$ denotes the initial condition, with
$V_{p}^{\delta_{u}}=\left(\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}),\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\right)_{1-\frac{1}{p},p}.$
Note that here, and in the remainder of the section, we drop the dependence of $u$ on $\varepsilon$.
In Section 4.1, we investigate the properties of the weak Stokes operator $A_{w}$, in particular the $\mathcal{H}^{\infty}$-calculus and maximal regularity. Then, in Section 4.2 we apply these results to prove a global for small-data result for (4.1), according to the mild notion of solution stated in Definition 2.4.
4.1. Stokes operator in weak setting, $\mathcal{H}^{\infty}$-calculus and maximal regularity
We start this section by recalling how the weak Stokes operator $A_{w}$ inherits the property of a bounded $\mathcal{H}^{\infty}$-calculus with $\mathcal{H}^{\infty}$-angle $\varphi^{\infty}_{A_{w}}=0$ from the (strong) Stokes operator $A$. This is obtained by following the approach in [25, Section 5].
From this, it follows the maximal regularity property for the (weak) Stokes operator and the mixed derivative theorem, which are summarised in Proposition 4.2 and Lemma 4.3, respectively. Then, we conclude the section showing how it is possible to estimate the non-linear convective term $u\cdot\nabla u$ in $L^{p}(0,T;\prescript{}{0}{}H^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))$ using the maximal regularity space $\mathbb{E}_{T,p}^{\delta_{u}}$, for any $p>\frac{2}{1-\delta_{u}}$. This is key to prove the global existence result for small data in Section 4.2.
Proposition 4.1.
The weak Stokes operator
$A_{w}\colon\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\to\prescript{}{0}{H^{-\frac{1}{2}-\delta_{u}}_{\sigma}}(\mathcal{D})$
admits a bounded $\mathcal{H}^{\infty}$-calculus on $\prescript{}{0}{}H^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})$ with angle $\varphi^{\infty}_{A_{w}}=0$.
Proof.
Let $A_{0}=A$ and $X_{0}=L^{2}_{\sigma}(\mathcal{D})$. By [5, Theorem V.1.5.1 and Theorem 1.5.4], it follows that the pair $(X_{0},A_{0})$ generates an interpolation-extrapolation scale $(X_{\alpha},A_{\alpha})$ with respect to the complex interpolation functor. Note that for $\alpha\in(0,1)$, $A_{\alpha}$ is the $X_{\alpha}$-realization of $A_{0}$ (the restriction of $A_{0}$ to $X_{\alpha}$) and
$X_{\alpha}=D(A_{0}^{\alpha}).$
Let $X_{0}^{\sharp}:=(X_{0})^{\prime}$ and $A_{0}^{\sharp}:=(A_{0})^{\prime}$ with $D(A_{0}^{\sharp})=:X_{1}^{\sharp}$. Then $(X_{0}^{\sharp},A_{0}^{\sharp})$ generates an interpolation-extrapolation scale $(X_{\alpha}^{\sharp},A_{\alpha}^{\sharp})$, the dual scale, and by [5, Theorem V.1.5.12], it holds that
$(X_{\alpha})^{\prime}=X^{\sharp}_{-\alpha}\quad\text{and}\quad(A_{\alpha})^{\prime}=A^{\sharp}_{-\alpha}$
for $\alpha\in\mathbb{R}$. Further, the domain of the fractional power of the Stokes operator can be characterised as
$X_{\alpha}=D((-\Delta)^{\alpha})\cap L^{2}_{\sigma}(\mathcal{D}),\quad\alpha\in(0,1),$
see [15]. In the particular case of
$\alpha=1-\theta=\frac{3}{4}-\frac{\delta_{u}}{2},$
then from the characterisation of the fractional powers of the Dirichlet Laplacian, we get
$X_{1-\theta}=\left\{u\in H^{\frac{3}{2}-\delta_{u}}(\mathcal{D})\ :\ u|_{\partial\mathcal{D}}=0\right\}\cap L^{2}_{\sigma}(\mathcal{D}),$
where $u|_{\partial\mathcal{D}}$ is understood in trace-sense. Hence, the operator $A_{-\theta}$ from the scale $(X_{\alpha},A_{\alpha})$, $\alpha\in\mathbb{R},$ is given by
$A_{-\theta}\colon X_{1-\theta}\to X_{-\theta},$
where by reflexivity $X_{-\theta}=\left(X_{\theta}^{\sharp}\right)^{\prime}$. Since $A_{0}^{\sharp}\in\mathcal{H}^{\infty}(X_{0}^{\sharp})$, we have
$X^{\sharp}_{\theta}=D((A_{0}^{\sharp})^{\theta})=[X_{0}^{\sharp},X_{1}^{\sharp}]_{\theta}=\left\{u\in H^{\frac{1}{2}+\delta_{u}}(\mathcal{D})\ :\ u|_{\partial\mathcal{D}}=0\right\}\cap L^{2}_{\sigma}(\mathcal{D}).$
Moreover, we have $A_{-\theta}=(A^{\sharp}_{\theta})^{\prime}$. Since $A_{0}^{\sharp}$ admits a bounded $\mathcal{H}^{\infty}$-calculus on $X_{0}^{\sharp}$ with angle 0, by duality the operator $A_{-\theta}\colon X_{1-\theta}\to X_{-\theta}$ has a bounded $\mathcal{H}^{\infty}$-calculus with $\mathcal{H}^{\infty}$-angle $\varphi^{\infty}_{A_{-\theta}}=0.$
It remains to identify $A_{-\theta}$ with the weak Stokes operator $A_{w}$. Since $A_{-\theta}$ is the closure of $A_{0}$ in $X_{-\theta}$, it coincides with $A_{0}$ on the dense subspace $X_{1}$. For any $u\in X_{1}$ and $v\in X_{1}^{\sharp}$, we have via integration by parts
$\langle A_{-\theta}u,v\rangle=(A_{0}u,v)_{L^{2}(\mathcal{D})}=\int_{\mathcal{D}}\nabla u\cdot\nabla v\,dx.$
Since $X_{1}^{\sharp}$ is dense in $X_{\theta}^{\sharp}$, this identity extends to all $v\in X_{\theta}^{\sharp}$ by interpreting the integral as duality pairing. Note that for $u\in X_{1-\theta}$ and $v\in X_{\theta}^{\sharp}$, the gradients satisfy $\nabla u\in H^{\frac{1}{2}-\delta_{u}}(\mathcal{D})$ and $\nabla v\in H^{-\frac{1}{2}+\delta_{u}}(\mathcal{D})$. Thus, the identity extends continuously with respect to $u$ to the duality pairing
$\langle A_{-\theta}u,v\rangle=\langle\nabla u,\nabla v\rangle_{{H^{\frac{1}{2}-\delta_{u}}},{H^{-\frac{1}{2}+\delta_{u}}}}.$
Since $X_{1}$ is dense in $X_{1-\theta}$, this identity holds for all $u\in X_{1-\theta}=\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})$ and $v\in X_{\theta}^{\sharp}=\prescript{}{0}{H}^{\frac{1}{2}+\delta_{u}}_{\sigma}(\mathcal{D})$. Therefore, the abstract operator $A_{-\theta}$ coincides with the weak Stokes operator $A_{w}$ defined in (2.1), and $A_{w}$ inherits the bounded $\mathcal{H}^{\infty}$-calculus. ∎
The first result that we recall is the maximal regularity of the weak Stokes operator, which is a consequence of the $\mathcal{H}^{\infty}$-calculus recalled above, see [24, Section II.4].
Proposition 4.2 (Maximal regularity for the weak-Stokes operator).
Let $T>0$ be a given time, $J=(0,T)$, $\delta_{u}\in(0,1)$. Consider the problem
[…]
  1. (i)
    If $g\in L^{p}(J;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))$ and $z_{0}\in V_{p}^{\delta_{u}}$, then there exists a unique solution
    $z\in\mathbb{E}_{T,p}^{\delta_{u}}=W^{1,p}(J;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))\cap L^{p}(J;\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))$
    of (4.2).
  2. (ii)
    There exists $C=C(p,\delta_{u},T)>0$ such that
    (4.3) $\norm{z}_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq C\left(\norm{g}_{L^{p}(J;H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}+\norm{z_{0}}_{V_{p}^{\delta_{u}}}\right).$
Second, we recall the mixed derivative theorem, see e.g. [24, Corollary 4.5.10].
Lemma 4.3 (Mixed derivative).
Let $T>0$ be a given time and $J=(0,T)$. Then
(4.4) $\mathbb{E}_{T,p}^{\delta_{u}}=W^{1,p}(J;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))\cap L^{p}(J;\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))\hookrightarrow H^{r,p}(J;D(A_{w}^{1-r})),$
for any $r\in[0,1].$
We conclude with an auxiliary result, which is based on Lemma 4.3, for the estimate of the non-linear convective term of the NSEs (4.1).
Lemma 4.4.
Let $T>0$ be a given time, $\delta_{u}\in(0,1)$ and set $J=(0,T)$. If $p>\frac{2}{1-\delta_{u}}$, then there exists $C=C(p,\delta_{u},T)>0$ such that
$\norm{P(u\cdot\nabla v)}_{L^{p}(J;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))}\leq C\norm{u}_{\mathbb{E}_{T,p}^{\delta_{u}}}\norm{v}_{\mathbb{E}_{T,p}^{\delta_{u}}}$
for any $u,v\in\mathbb{E}_{T,p}^{\delta_{u}}$, where $P:H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})\to\prescript{}{0}{H^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})}$ is the extension of the Helmholtz projection.
Proof.
Assume that $s_{1},s_{2}\in\mathbb{R}$ are such that the following pointwise multiplication is bilinear and continuous
(4.5) $W^{s_{1},2}(\mathcal{D})\times W^{s_{2},2}(\mathcal{D})$ $\hookrightarrow W^{-\frac{1}{2}-\delta_{u},2}(\mathcal{D})$
$(f,g)$ $\mapsto fg.$
Combining the previous multiplication embedding with the boundedness of $P$ on $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$ and the norm equivalence $H^{s}(\mathcal{D})\cong W^{s,2}(\mathcal{D})$, there exists $C=C(s_{1},s_{2},\delta_{u})>0$ such that
$\norm{P(u\cdot\nabla v)}_{\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})}\leq C\norm{u}_{W^{s_{1},2}(\mathcal{D})}\norm{\nabla v}_{W^{s_{2},2}(\mathcal{D})}\leq C\norm{u}_{W^{s_{1},2}(\mathcal{D})}\norm{v}_{W^{1+s_{2},2}(\mathcal{D})}.$
Considering the $L^{p}$ norm in time and applying Hölder’s inequality, we obtain
$\begin{split}\norm{P(u\cdot\nabla v)}_{L^{p}(J;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))}&\leq C\norm{u\cdot\nabla v}_{L^{p}(J;H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}\\ &\leq C\norm{u}_{L^{\infty}(J;W^{s_{1},2}(\mathcal{D}))}\norm{v}_{L^{p}(J;W^{1+s_{2},2}(\mathcal{D}))}.\end{split}$
Assuming that
(4.6) $\mathbb{E}_{T,p}^{\delta_{u}}\hookrightarrow L^{\infty}(J;H^{s_{1}}(\mathcal{D}))$
and
(4.7) $\mathbb{E}_{T,p}^{\delta_{u}}\hookrightarrow L^{p}(J;H^{1+s_{2}}(\mathcal{D})),$
we conclude, from the norm equivalence $H^{s}(\mathcal{D})\cong W^{s,2}(\mathcal{D})$, that
$\norm{P(u\cdot\nabla v)}_{L^{p}(J;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))}\leq C\norm{u}_{\mathbb{E}_{T,p}^{\delta_{u}}}\norm{v}_{\mathbb{E}_{T,p}^{\delta_{u}}}.$
It remains to check that (4.5), (4.6) and (4.7) hold. First, using Lemma 4.3, it holds
$\mathbb{E}_{T,p}^{\delta_{u}}\hookrightarrow H^{r,p}(J;D(A_{w}^{1-r})),$
for any $r\in[0,1]$. Second, using the notation of the proof of Proposition 4.1, recall that the pair $(X_{0},A_{0})$, where $A_{0}$ denotes the Stokes operator $A_{0}=A$, and $X_{0}=L^{2}_{\sigma}(\mathcal{D})$, generates the interpolation-extrapolation scale $(X_{\alpha},A_{\alpha})$. Further, the weak Stokes operator is $A_{w}=A_{-\theta}:X_{1-\theta}\to X_{-\theta}$ and $D(A_{w}^{1-r})=X_{-\theta+(1-r)}=\prescript{}{0}{H}_{\sigma}^{2(1-\theta-r)}(\mathcal{D})$. Since $1-\theta=\frac{3}{4}-\frac{\delta_{u}}{2}$ (see the proof of Proposition 4.1), we have $D(A_{w}^{1-r})=\prescript{}{0}{H}_{\sigma}^{\frac{3}{2}-\delta_{u}-2r}(\mathcal{D})$, and in conclusion
$\mathbb{E}_{T,p}^{\delta_{u}}\hookrightarrow H^{r,p}(J;D(A_{w}^{1-r}))\hookrightarrow H^{r,p}\bigl(J;H^{\frac{3}{2}-\delta_{u}-2r}(\mathcal{D})\bigr).$
Further, it can be checked that (4.6) holds if
(4.8) $\left\{\begin{aligned} r&>\frac{1}{p},\\ \frac{3}{2}-\delta_{u}-2r&\geq s_{1},\end{aligned}\right$.
is satisfied for some $r\in[0,1]$. On the other hand, (4.7) holds if $s_{2}\leq\frac{1}{2}-\delta_{u}$. Lastly, choosing
$s_{1}:=\frac{3}{2}-\delta_{u}-2r,\quad s_{2}:=\frac{1}{2}-\delta_{u},$
the pointwise multiplication embedding (4.5) holds, thanks to [6, Theorem 8.2] if $\delta_{u}\in(0,\frac{1}{2})$ and [6, Theorem 8.1] if $\delta_{u}\in[\frac{1}{2},1)$, if $r<\frac{1-\delta_{u}}{2}$. Thus, we can find $r\in(0,1)$ such that (4.5) and (4.8) hold if
$\frac{1}{p}<\frac{1-\delta_{u}}{2},$
which is true thanks to our hypothesis $p>\frac{2}{1-\delta_{u}}$. This completes the proof. ∎
4.2. Global for small data for the 3D Navier-Stokes problem
The auxiliary results presented in the previous section will be employed to demonstrate Theorem 2.5, i.e. the global existence result for small data for the 3D Navier-Stokes problem (4.1), within the weak setting for the Stokes operator $A_{w}$. In particular, if we assume that the initial condition $u_{0}$ and the forcing term $f^{\theta}$ are sufficiently small, we show that the solution, in the sense of Definition 2.4, exists on the whole time interval $[0,T]$.
Proof of Theorem 2.5.
Let $v_{*}$ denote the reference solution for (4.1), i.e., the solution to the linear problem
(4.9) $\left\{\begin{aligned} \partial_{t}v_{*}+A_{w}v_{*}&=Pf^{\theta},&\quad&t\in(0,T),\\ v_{*}|_{t=0}&=u_{0},\end{aligned}\right$.
and define the ball
$\mathbb{B}\left(v_{*},\frac{\eta}{2}\right)=\left\{v^{\prime}\in\mathbb{E}_{T,p}^{\delta_{u}}\;:\;\norm{v^{\prime}-v_{*}}_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq\frac{\eta}{2}\right\}\subset\mathbb{E}_{T,p}^{\delta_{u}}.$
Consider the map
$\Gamma\colon\mathbb{B}\left(v_{*},\frac{\eta}{2}\right)\subset\mathbb{E}_{T,p}^{\delta_{u}}$ $\to\mathbb{E}_{T,p}^{\delta_{u}}$
$\Gamma(v^{\prime})$ $=v,$
where $v$ is the unique solution of
[…]
  1. (i)
    $\norm{\Gamma(v^{\prime})-v_{*}}_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq\frac{\eta}{2}$ for all $v^{\prime}\in\mathbb{B}(v_{*},\frac{\eta}{2})$,
  2. (ii)
    $\norm{\Gamma(v_{1}^{\prime})-\Gamma(v_{2}^{\prime})}_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq d\cdot\norm{v_{1}^{\prime}-v_{2}^{\prime}}_{\mathbb{E}_{T,p}^{\delta_{u}}}$, where $d\in[0,1)$, for all $v_{1}^{\prime},v_{2}^{\prime}\in\mathbb{B}(v_{*},\frac{\eta}{2}).$
(i). Let $v=\Gamma(v^{\prime})$ and define $w:=v-v_{*}$, which satisfies
(4.10) $\left\{\begin{aligned} \partial_{t}w+A_{w}w&=P(-v^{\prime}\cdot\nabla v^{\prime}),&\quad&t\in(0,T),\\ w|_{t=0}&=0.\end{aligned}\right$.
By the maximal regularity of the weak Stokes operator (see Proposition 4.2), Lemma 4.4 and Jensen’s inequality, we obtain
$\begin{split}\norm{w}_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq C\norm{P(v^{\prime}\cdot\nabla v^{\prime})}_{L^{p}(J;\prescript{}{0}{H}^{-\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D}))}\leq C\norm{v^{\prime}}_{\mathbb{E}_{T,p}^{\delta_{u}}}^{2}\leq 2C\left(\norm{v^{\prime}-v_{*}}^{2}_{\mathbb{E}_{T,p}^{\delta_{u}}}+\norm{v_{*}}_{\mathbb{E}_{T,p}^{\delta_{u}}}^{2}\right),\end{split}$
for a positive constant $C=C(p,\delta_{u},T)$ and $J=(0,T).$ Further, again by the maximal regularity of the weak Stokes operator, there exists a constant $M=M(p,\delta_{u},T)$ such that
(4.11) $\begin{split}\norm{v_{*}}_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq M\left(\norm{f^{\theta}}_{L^{p}(J;H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}+\norm{u_{0}}_{V_{p}^{\delta_{u}}}\right)\leq\frac{\eta}{2}.\end{split}$
Thus, we have
(4.12) $\begin{split}\norm{v-v_{*}}_{\mathbb{E}_{T,p}^{\delta_{u}}}=\norm{w}_{\mathbb{E}_{T,p}^{\delta_{u}}}&\leq 2C\left(\norm{v^{\prime}-v_{*}}^{2}_{\mathbb{E}_{T,p}^{\delta_{u}}}+\norm{v_{*}}_{\mathbb{E}_{T,p}^{\delta_{u}}}^{2}\right)\leq 2C\left(\frac{\eta^{2}}{4}+\frac{\eta^{2}}{4}\right)=C\eta^{2}.\end{split}$
By choosing $\eta$ such that
$\eta^{2}\leq\frac{\eta}{2C},$
we conclude that $\Gamma$ maps $\mathbb{B}\left(v_{*},\frac{\eta}{2}\right)$ into itself.
(ii). Let $v_{i}=\Gamma(v_{i}^{\prime})$ for $i=1,2,$ and define $\tilde{w}:=v_{1}-v_{2}$, $\tilde{w}^{\prime}=v_{1}^{\prime}-v_{2}^{\prime}$. Then, $\tilde{w}$ satisfies
(4.13) $\left\{\begin{aligned} \partial_{t}\tilde{w}+A_{w}\tilde{w}&=P(-v_{1}^{\prime}\cdot\nabla v_{1}^{\prime}+v_{2}^{\prime}\cdot\nabla v_{2}^{\prime}),&t\in(0,T),\\ \tilde{w}|_{t=0}&=0.\end{aligned}\right$.
By the maximal regularity of the Stokes operator, Lemma 4.4, the definition of $v_{i}^{\prime}$ and (4.11), we have
$\begin{split}\norm{\Gamma(v_{1}^{\prime})-\Gamma(v_{2}^{\prime})}_{\mathbb{E}_{T,p}^{\delta_{u}}}&\leq C\norm{-\tilde{w}^{\prime}\cdot\nabla v_{1}^{\prime}-v_{2}^{\prime}\cdot\nabla\tilde{w}^{\prime}}_{L^{p}(J;H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}\\ &\leq C\norm{\tilde{w}^{\prime}}_{\mathbb{E}_{T,p}^{\delta_{u}}}\left(\norm{v_{1}^{\prime}}_{\mathbb{E}_{T,p}^{\delta_{u}}}+\norm{v_{2}^{\prime}}_{\mathbb{E}_{T,p}^{\delta_{u}}}\right)\\ &\leq C\norm{\tilde{w}^{\prime}}_{\mathbb{E}_{T,p}^{\delta_{u}}}\left(\eta+\eta\right)\\ &=2C\norm{v_{1}^{\prime}-v_{2}^{\prime}}_{\mathbb{E}_{T,p}^{\delta_{u}}}\eta,\end{split}$
where $C=C(p,\delta_{u},T)$ is a positive constant changing from line to line. Up to choosing $\eta$ sufficiently small, (ii) is verified.
In conclusion, by the Banach fixed point theorem, there exists a unique fixed point $v=\Gamma(v)\in\mathbb{B}(v_{*},\frac{\eta}{2}).$ Furthermore, by the triangle inequality, (4.11), and (4.12), we also obtain
$\norm{v}_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq\norm{v-v_{*}}_{\mathbb{E}_{T,p}^{\delta_{u}}}+\norm{v_{*}}_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq\frac{\eta}{2}+\frac{\eta}{2}=\eta.$
This concludes the proof. ∎
5. Well-posedness for the 3D-coupled temperature-velocity model with boundary noise
We are now in position to prove the main result of our work.
Proof of Theorem 2.8.
Fix $T>0$, $s\in[0,\frac{1}{2})$ and $p>4$. We proceed to verify that parameters satisfying the hypotheses exist. First, we choose $\gamma=\frac{1}{4}+\varepsilon_{\gamma}$ with $\varepsilon_{\gamma}>0$ small. Recall $a(s,\gamma)=\frac{3}{4}+\frac{s}{2}+\varepsilon_{\gamma}$. Since $s<1/2$, we have $a(s,\gamma)<1$ for small $\varepsilon_{\gamma}$. We require $\lambda_{\max}>\lambda_{\min}$, i.e., $p(1-a(s,\gamma))>\frac{1}{2}-s$. As $\varepsilon_{\gamma}\to 0$, this condition becomes
$p\left(1-\frac{3}{4}-\frac{s}{2}\right)>\frac{1}{2}-s\iff p\frac{1-2s}{4}>\frac{1-2s}{2}\iff p>2.$
Since we assumed $p>4>2$, the previous condition is satisfied. Next, we choose $\lambda\in(\lambda_{\min},\lambda_{\max})\cap(0,1)$ and define $\overline{\delta}_{\theta},\overline{\delta}_{u}$ as in the statement. We must show that the interval for $\delta_{u}$ is non-empty, i.e.,
$\max\{\delta_{\theta},\tfrac{1}{2}-s\}\leq\delta_{u}<\min\{\overline{\delta}_{u},1-\tfrac{2}{p}\}.$
We can always choose $\delta_{\theta}$ small enough such that $\delta_{\theta}\leq 1/2-s$. Thus, we need to ensure $\frac{1}{2}-s<\min\{\overline{\delta}_{u},1-\frac{2}{p}\}$. This is verified, up to choosing $\varepsilon_{\gamma}>0$ small enough, with the same arguments as in Remark 2.9. Finally, note that the condition $\delta_{u}<1-2/p$ implies $p>\frac{2}{1-\delta_{u}}$, which ensures the maximal regularity applicability in Theorem 2.5.
Define
$\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2},\qquad\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4}.$
Let $\tilde{\eta}=\tilde{\eta}(p,\delta_{u},T)>0$ and $M=M(p,\delta_{u},T)>0$ be the constants given by Theorem 2.5. Since $\delta_{u}\geq\max\{\delta_{\theta},\frac{1}{2}-s\}$, the embeddings
$H^{-2\alpha_{\theta}}(\mathcal{D})\hookrightarrow H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}),\quad W^{s,\frac{6}{5}}(\mathcal{D})\hookrightarrow H^{s-1}(\mathcal{D})\hookrightarrow H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}),$
hold. Thus, there exists a constant $C_{emb}=C_{emb}(\delta_{\theta},\delta_{u},s)$ such that for all $g_{1}\in H^{-2\alpha_{\theta}}(\mathcal{D})$ and $g_{2}\in W^{s,\frac{6}{5}}(\mathcal{D})$, we have
$\|g_{1}\|_{H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})}\leq C_{emb}\|g_{1}\|_{H^{-2\alpha_{\theta}}(\mathcal{D})},\quad\|g_{2}\|_{H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})}\leq C_{emb}\|g_{2}\|_{W^{s,\frac{6}{5}}(\mathcal{D})}.$
[…]
$\tau^{\varepsilon}:=\inf\left\{t>0\ :\ \norm{Z^{\varepsilon}_{t}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}>\frac{\eta}{8\widetilde{M}}\right\}$,
with the understanding that $\tau^{\varepsilon}=T$ if the previous set is empty. Note that
$\left\{\omega\ :\ \tau^{\varepsilon}=T\right\}=\left\{\omega\ :\ \sup_{0\leq t\leq T}\norm{Z^{\varepsilon}_{t}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}\leq\frac{\eta}{8\widetilde{M}}\right\}$.
From Proposition 2.1, we deduce that
$\mathbb{P}\left(\tau^{\varepsilon}=T\right)\geq 1-\frac{64\widetilde{M}^{2}\varepsilon}{\eta^{2}}\,C(\delta_{\theta},T)\,\sum\limits_{k}\lambda_{k}^{2}\norm{(-\Delta)^{\beta_{\theta}}De_{k}}_{2}^{2},$
where $C=C(\delta_{\theta},T)$ is a positive constant. Consider
$\mathbb{B}_{1}:=\left\{z\in C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}}(\mathcal{D}))\ :\ \norm{z}_{C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}}(\mathcal{D}))}\leq\frac{\eta}{8\widetilde{M}}\right\}$,
and the map
$\begin{split}\Gamma\colon\mathbb{B}_{1}&\to C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}}(\mathcal{D}))\\ \zeta^{\prime}&\mapsto\zeta^{\varepsilon}\end{split}$
where
$\zeta_{t}^{\varepsilon}:=e^{t\Delta}\theta_{0}-\int_{0}^{t}e^{(t-r)\Delta}\left(u^{\varepsilon}_{r}\cdot\nabla\zeta^{\prime}_{r}\right)\,dr-\int_{0}^{t}e^{(t-r)\Delta}\left(u_{r}^{\varepsilon}\cdot\nabla Z^{\varepsilon}_{r}\right)\,dr,\quad 0\leq t\leq\tau^{\varepsilon}.\\$
Here $u^{\varepsilon}\in\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}$ is the unique solution of
$\left\{\begin{aligned} \partial_{t}u^{\varepsilon}+A_{w}u^{\varepsilon}&=P(-u^{\varepsilon}\cdot\nabla u^{\varepsilon}-\theta_{t}^{\varepsilon,\prime}e_{3}),\quad&t\in(0,\tau^{\varepsilon}),\\ u^{\varepsilon}|_{t=0}&=u_{0},\end{aligned}\right$.
where
$\theta_{t}^{\varepsilon,\prime}:=Z^{\varepsilon}_{t\wedge\tau^{\varepsilon}}+\zeta^{\prime}_{t}.$
Note that $u^{\varepsilon}$ is well defined thanks to Theorem 2.5, once we check the corresponding smallness condition on the forcing $\theta^{\varepsilon,\prime}e_{3}$. First, by the definition of $\tau^{\varepsilon}$ we have
$\sup_{0\leq t\leq T}\bigl\|Z^{\varepsilon}_{t\wedge\tau^{\varepsilon}}\bigr\|_{H^{-2\alpha_{\theta}}(\mathcal{D})}=\sup_{0\leq t\leq T}\bigl\|Z^{\varepsilon}_{t\wedge\tau^{\varepsilon}}\bigr\|_{H^{-\frac{1}{2}-\delta_{\theta}}(\mathcal{D})}\leq\frac{\eta}{8\widetilde{M}}$.
Further, by definition of $\mathbb{B}_{1}$, we also have
$\sup_{0\leq t\leq\tau^{\varepsilon}}\|\zeta_{t}^{\prime}\|_{W^{s,\frac{6}{5}}(\mathcal{D})}\leq\frac{\eta}{8\widetilde{M}}$.
Using the embedding constant $C_{emb}$ defined above, we estimate the forcing in $L^{p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))$
[…]
Since we defined $\widetilde{M}\geq MT^{1/p}C_{emb},$ we deduce
$\|\theta^{\varepsilon,\prime}\|_{L^{p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}\leq\frac{\eta}{4M}.$
This is exactly the smallness assumption required in Theorem 2.5, so $u^{\varepsilon}$ is well defined.
To apply the Banach fixed point theorem, we first check that $\Gamma$ maps $\mathbb{B}_{1}$ into itself. We can use (2.8) by Corollary 2.3 since its hypothesis, namely $\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $\delta_{u}<\overline{\delta}_{u}$ are satisfied; further, we have the embedding $\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}\hookrightarrow L^{\frac{p}{\lambda+\delta_{u}}}(0,\tau^{\varepsilon};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})),$ which holds for any $0<\lambda\leq 1-\delta_{u}$. Thus,
[…]
Thus, if $\eta$ is small enough, $\Gamma(\mathbb{B}_{1})\subset\mathbb{B}_{1}$.
Now we verify the contraction property. Consider $\zeta_{i}^{\varepsilon}=\Gamma(\zeta_{i}^{\prime})$, $i=1,2$. By the definition of $\Gamma$, for every $t\in[0,\tau^{\varepsilon}]$ we have
$\begin{split}\zeta_{1}^{\varepsilon}(t)-\zeta_{2}^{\varepsilon}(t)&=-\int_{0}^{t}e^{(t-r)\Delta}\Big(u_{1,r}^{\varepsilon}\cdot\nabla\zeta_{1,r}^{\prime}-u_{2,r}^{\varepsilon}\cdot\nabla\zeta_{2,r}^{\prime}\Big)\,dr\\ &\quad-\int_{0}^{t}e^{(t-r)\Delta}\Big(u_{1,r}^{\varepsilon}-u_{2,r}^{\varepsilon}\Big)\cdot\nabla Z_{r}^{\varepsilon}\,dr.\end{split}$
We split the first integral by adding and subtracting the same quantity and then, using Corollary 3.3 and the bilinear estimate proved in (3.29), we obtain
$\begin{split}\norm{\zeta_{1}^{\varepsilon}-\zeta_{2}^{\varepsilon}}_{C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}})}&\leq C\biggl(\|u^{\varepsilon}_{1}\|_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}\|\zeta_{1}^{\prime}-\zeta_{2}^{\prime}\|_{C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}})}\\ &\quad+\|u^{\varepsilon}_{1}-u^{\varepsilon}_{2}\|_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}\|\zeta_{2}^{\prime}\|_{C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}})}\\ &\quad+\|u^{\varepsilon}_{1}-u^{\varepsilon}_{2}\|_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}\|Z^{\varepsilon}\|_{C(0,\tau^{\varepsilon};H^{-2\alpha_{\theta}})}\biggr).\end{split}$
Let $w^{\varepsilon}:=u_{1}^{\varepsilon}-u_{2}^{\varepsilon}$. It satisfies
$\left\{\begin{aligned} \partial_{t}w^{\varepsilon}+A_{w}w^{\varepsilon}&=P(-u_{1}^{\varepsilon}\cdot\nabla u^{\varepsilon}_{1}+u_{2}^{\varepsilon}\cdot\nabla u^{\varepsilon}_{2}-\left(\zeta_{t,1}^{\prime}-\zeta_{t,2}^{\prime}\right)e_{3}),\quad&t\in(0,\tau^{\varepsilon}),\\ w^{\varepsilon}|_{t=0}&=0.\end{aligned}\right$.
By the maximal regularity of the weak Stokes operator (Proposition 4.2) and Lemma 4.4, there exists $C=C(p,\delta_{u},T)$ such that
$\begin{split}\norm{w^{\varepsilon}}_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}&\leq C\left(\norm{w^{\varepsilon}\cdot\nabla u_{1}^{\varepsilon}}_{L^{p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}+\norm{u_{2}^{\varepsilon}\cdot\nabla w^{\varepsilon}}_{L^{p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}\right.\\ &\quad\left.+\norm{\zeta^{\prime}_{1}-\zeta^{\prime}_{2}}_{L^{p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}\right)\\ &\leq C\left(\norm{w^{\varepsilon}}_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}\norm{u_{1}^{\varepsilon}}_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}+\norm{u_{2}^{\varepsilon}}_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}\norm{w^{\varepsilon}}_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}+\norm{\zeta^{\prime}_{1}-\zeta^{\prime}_{2}}_{L^{p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}\right)\\ &\leq C\left(2\eta\norm{w^{\varepsilon}}_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}+\norm{\zeta^{\prime}_{1}-\zeta^{\prime}_{2}}_{L^{p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}\right).\end{split}$
Up to choosing $\eta$ sufficiently small, using the embedding $W^{s,\frac{6}{5}}(\mathcal{D})\hookrightarrow H^{s-1}(\mathcal{D})\hookrightarrow H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$, which holds for $\delta_{u}\geq\frac{1}{2}-s,$ we obtain
$\norm{w^{\varepsilon}}_{\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}}\leq C\norm{\zeta^{\prime}_{1}-\zeta^{\prime}_{2}}_{L^{p}(0,\tau^{\varepsilon};H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}\leq C\norm{\zeta^{\prime}_{1}-\zeta^{\prime}_{2}}_{C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}}(\mathcal{D}))}.$
Up to renaming $C$, we get
$\begin{split}\norm{\zeta_{1}^{\varepsilon}-\zeta_{2}^{\varepsilon}}_{C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}}(\mathcal{D}))}&\leq C\left(\eta\norm{\zeta_{1}^{\prime}-\zeta_{2}^{\prime}}_{C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}}(\mathcal{D}))}+\norm{\zeta_{1}^{\prime}-\zeta_{2}^{\prime}}_{C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}}(\mathcal{D}))}\frac{\eta}{8\widetilde{M}}\right)\\ &=d\norm{\zeta_{1}^{\prime}-\zeta_{2}^{\prime}}_{C(0,\tau^{\varepsilon};W^{s,\frac{6}{5}}(\mathcal{D}))}\end{split}$
[…]
Lastly, regarding the adaptedness of the solution process, fix a time $t_{0}\in[0,T]$. For any $\omega\in\Omega$, the contraction principle applied in the proof above, starting with a deterministic initial condition, generates an adapted sequence $(\tilde{u}^{\varepsilon,t_{0}}_{n}(\omega),\tilde{\theta}^{\varepsilon,t_{0}}_{n}(\omega))_{n}$ which converges to a limiting sequence $(\tilde{u}^{\varepsilon,t_{0}}(\omega),\tilde{\theta}^{\varepsilon,t_{0}}(\omega))$. This latter stochastic process is adapted, being the limit of adapted processes. By uniqueness, the restriction to $0,t_{0}]$ of the solution $(u^{\varepsilon}(\omega),\tilde{\theta}^{\varepsilon}(\omega))$ needs to coincide with the adapted solution $(\tilde{u}^{\varepsilon,t_{0}},\tilde{\theta}^{\varepsilon,t_{0}})$. This proves adaptedness. The progressive measurability of the solution can be obtained by a similar reasoning. The proof is thus complete. ∎
Acknowledgments
G.D.S. acknowledges the support from the DFG project FOR 5528. M.L. is supported by the the Italian national inter-university PhD course in Sustainable Development and Climate change. M.L. produced this work while attending the PhD programme in PhD in Sustainable Development And Climate Change at the University School for Advanced Studies IUSS Pavia, Cycle XXXVIII, with the support of a scholarship financed by the Ministerial Decree no. 351 of 9th April 2022, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 "Education and Research", Component 1 "Enhancement of the offer of educational services: from nurseries to universities” - Investment 3.4 “Advanced teaching and university skills”. The authors thank Franco Flandoli and Eliseo Luongo for the fruitful discussion on the topic.

### v3

@@
 The parameter $\varepsilon$ influences the probability of global existence up to time $T$.
-Indeed, the three-dimensional (3D) Navier-Stokes equations are only known to be globally well‑posed under smallness conditions on initial data and forcing (see, e.g., [8, Chapter 9]).
+Indeed, the three-dimensional (3D) Navier-Stokes equations are only known to be globally well‑posed under smallness conditions on initial data and forcing (see, e.g., [9, Chapter 9]).
 Since stochastic boundary forcing enters the velocity equation through buoyancy coupling, we impose a small‑noise regime in order to retain a nonzero probability of well‑posedness on $[0,T]$.
@@
 $\delta_{u}\geq\max\{\delta_{\theta},\ \tfrac{1}{2}-s\},$
-where $s\in[0,1/2)$ is the Sobolev regularity of the initial temperature $\theta_{0}\in W^{s,6/5}(\mathcal{D})$.
+where $s\in(0,1/2)$ is the Sobolev regularity of the initial temperature $\theta_{0}\in W^{s,6/5}(\mathcal{D})$.
 Under this condition (and for $p>4$ in the maximal-regularity framework), we prove the existence and uniqueness of the coupled system on the interval $0,\tau^{\varepsilon}]$, satisfying
@@
 Furthermore, we establish the high-probability estimate $\mathbb{P}(\tau^{\varepsilon}=T)\geq 1-C\varepsilon$.
-As $s\uparrow 1/2$, the constraint from the initial datum disappears and one may choose $\delta_{u}=\delta_{\theta}$, recovering a symmetric regularity regime.
 It is worth underlining two aspects of the boundary noise: its mathematical difficulties, and its physical meaning.
 Firstly, Da Prato and Zabczyk were the first to rigorously demonstrate in [10] that solutions to the heat equation subject to Dirichlet boundary noise exhibit significantly reduced regularity compared to solutions driven by interior stochastic forcing, even in one spatial dimension.
-Specifically, due to the presence of boundary noise, the best regularity that one can hope for, even in the linear case, is continuity in time with values in a negative-order Bessel potential space, namely $C_{t}(H_{x}^{-\frac{1}{2}-\delta_{\theta}})$ for any $\delta_{\theta}>0$.
+Specifically, due to the presence of boundary noise, the best regularity that one can hope for, even in the linear case, is continuity in time with values in a negative-order Bessel potential space, namely $C(0,T;H^{-\frac{1}{2}-\delta_{\theta}}(\mathcal{D}))$ for any $\delta_{\theta}>0$ and any final time $T>0$.
 This limitation reflects the difficulties of noisy Dirichlet boundary conditions.
@@
 This constitutes the optimal regularity according to the maximal regularity theory for the Stokes operator on $H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D})$, with suitable divergence-free and boundary conditions on the spaces.
-Our maximal regularity approach follows the framework developed by Prüss and Wilke in [25, Section 5].
-Secondly, from a physical point of view, stochastic boundary forcing encapsulates the influence of unresolved, rapid fluctuations, such as boundary-layer instabilities, or small-scale convection, that cannot be parametrized deterministically at the scale of our model.
-This philosophy traces back to Hasselmann’s stochastic climate paradigm [18], in which slow, large-scale dynamics are driven by fast, random perturbations.
+Our maximal regularity approach follows the framework developed by Prüss and Wilke in [27, Section 5].
+Secondly, from a physical point of view, stochastic boundary forcing represents the effect of unresolved fast fluctuations, such as boundary-layer instabilities or small-scale convection, that cannot be described deterministically at the scale of our model.
+Related approaches that model boundary-layer effects through non-standard boundary mechanisms can be found, for instance, in [7].
+In climate science, this philosophy traces back to Hasselmann’s stochastic climate paradigm [19], in which slow, large-scale dynamics are driven by fast, random perturbations.
 The boundary-noise framework thus provides a mathematically tractable way to “close” the system.
@@
 We denote by $(W_{t})_{t}$ a $\mathcal{Q}$-Wiener process on $L^{2}(\partial\mathcal{D})$ defined on $\left(\Omega,\mathcal{F},\left(\mathcal{F}_{t}\right)_{t},\mathbb{P}\right)$, and represented by
-$W(t,x)=\sum_{k\geq 0}\lambda_{k}e_{k}(x)\beta_{k}(t),\quad t\geq 0,\ x\in\partial\mathcal{D},$
+$W(t,x)=\sum_{k\geq 0}\lambda_{k}e_{k}(x)\beta_{k}(t),\quad t\geq 0,\ x\in\partial\mathcal{D}$
 where $(e_{k})_{k}\subset L^{2}(\partial\mathcal{D})$ is an orthonormal basis of $L^{2}(\partial\mathcal{D})$, $(\beta_{k}(t))_{k}$ are independent Brownian motions, and $(\lambda_{k})_{k}$ are the non-negative square roots of the eigenvalues of the covariance operator $\mathcal{Q}$.
@@
 $D((-\Delta)^{\alpha}):=\prescript{}{0}{H}^{2\alpha}(\mathcal{D}):=\left\{\begin{aligned} &H^{2\alpha}(\mathcal{D}),&\quad&0<\alpha<\frac{1}{4},\\ &H^{\frac{1}{2}}_{00}(\mathcal{D}),&\quad&\alpha=\frac{1}{4},\\ &H^{2\alpha}_{0}(\mathcal{D}),&\quad&\frac{1}{4}<\alpha\leq\frac{1}{2},\\ &H^{2\alpha}(\mathcal{D})\cap H^{1}_{0}(\mathcal{D}),&\quad&\frac{1}{2}<\alpha<1,\end{aligned}\right$.
-where $H^{2\alpha}_{00}(\mathcal{D})$ consists of all $u\in H^{\frac{1}{2}}(\mathcal{D})$ such that
+where $H^{\frac{1}{2}}_{00}(\mathcal{D})$ consists of all $u\in H^{\frac{1}{2}}(\mathcal{D})$ such that
 […]
@@
 $Au:=-P\Delta u,\quad D(A):=H^{2,2}(\mathcal{D})\cap H^{1,2}_{0}(\mathcal{D})\cap L^{2}_{\sigma}(\mathcal{D}).$
-In the Hilbert space $X_{0}=L^{2}_{\sigma}(\mathcal{D})$, the Stokes operator $A$ is a non-negative, self-adjoint operator with compact inverse; see, for instance, [27, Section 2].
+In the Hilbert space $X_{0}=L^{2}_{\sigma}(\mathcal{D})$, the Stokes operator $A$ is a non-negative, self-adjoint operator with compact inverse; see, for instance, [29, Section 2].
 In particular, $A$ is sectorial of angle 0 and A-A generates a bounded analytic $C_{0}$-semigroup on $X_{0}$; see, for example, [13, 20].
-Moreover, since $A$ is a non-negative self-adjoint operator on a Hilbert space, it admits a bounded $\mathcal{H}^{\infty}$-functional calculus of angle 0; see [17, Chapters 2 and 7].
+Moreover, since $A$ is a non-negative self-adjoint operator on a Hilbert space, it admits a bounded $\mathcal{H}^{\infty}$-functional calculus of angle 0; see [18, Chapters 2 and 7].
 In the specific case of the Stokes operator, the domains of the fractional powers $A^{\alpha}$ in $L^{2}_{\sigma}(\mathcal{D})$ have been identified in [16] and are given by
@@
 We conclude this section by describing the solution of the linear temperature problem associated to (2.2), i.e.
-by recalling the following result concerning the mild solution of (2.4), which follows from [9] and the factorisation trick for the stochastic convolution [11, Section 5.3.1].
+by recalling the following result concerning the mild solution of (2.4), which follows from [10] and the factorisation trick for the stochastic convolution [12, Section 5.3.1].
 Proposition 2.1.
@@
 2.3. Main results
-We are now in a position to state the main results of our work. The first one concerns the well-posedness of the system (2.5).
+We are now in a position to state the main results of our work. For this purpose, we begin by introducing some notation.
+Let $p>4$ and $s\in(0,\tfrac{1}{2})$. Choose
+(2.8) $\gamma:=\frac{1}{4}+\varepsilon_{\gamma},\qquad 0<\varepsilon_{\gamma}<\min\left\{\frac{1-2s}{4},\ \frac{(1-2s)(p-4)}{4p}\right\},$
+and set $a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}$ and $\lambda_{\max}:=p(1-a(s,\gamma))$. Then
+$a(s,\gamma)=\frac{3}{4}+\frac{s}{2}+\varepsilon_{\gamma}<1\quad\text{(since }\varepsilon_{\gamma}<\tfrac{1-2s}{4}\text{)},$
+and, using $1-a(s,\gamma)=\frac{1-2s}{4}-\varepsilon_{\gamma}$,
+(2.9) $\lambda_{\max}-\left(\frac{1}{2}-s\right)=p\left(\frac{1-2s}{4}-\varepsilon_{\gamma}\right)-\frac{1-2s}{2}=\frac{(1-2s)(p-2)}{4}-p\varepsilon_{\gamma}>0,$
+so $\lambda_{\max}>\frac{1}{2}-s$. Moreover,
+(2.10) $\lambda_{\max}-(1-2s)=p\left(\frac{1-2s}{4}-\varepsilon_{\gamma}\right)-(1-2s)=\frac{(1-2s)(p-4)}{4}-p\varepsilon_{\gamma}>0,$
+by (2.8). Hence the interval $(0,\min\{2s,\ \lambda_{\max}-(1-2s)\})$ is non-empty, and we may choose
+(2.11) $\lambda:=\left(\frac{1}{2}-s\right)+\varepsilon_{\lambda},\qquad 0<\varepsilon_{\lambda}<\min\left\{2s,\lambda_{\max}-(1-2s)\right\}.$
+In particular, $\lambda<\frac{1}{2}+s<1$ (since $\varepsilon_{\lambda}<2s$). Finally, define the thresholds
+(2.12) $\overline{\delta}_{\theta}:=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\},\qquad\overline{\delta}_{u}:=\min\left\{\lambda_{\max}-\lambda,1-\lambda\right\}.$
+The first main result concerns the well-posedness of the system (2.5)
 Theorem 2.2.
-Let $s\in[0,1/2)$ and $p>2$.
-Then there exist $\overline{\delta}_{\theta}=\overline{\delta}_{\theta}(s,p)>0$ and $\overline{\delta}_{u}=\overline{\delta}_{u}(s,p)>0$ such that for every $0<\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $0<\delta_{u}<\overline{\delta}_{u}$, the following holds.
-Let
-$\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2},\quad\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4},$
-and denote by $Z^{\varepsilon}_{t}=-\sqrt{\varepsilon}\int_{0}^{t}\Delta e^{(t-r)\Delta}D\,dW_{r}$ the continuous version in $H^{-2\alpha_{\theta}}(\mathcal{D})$ considered in Proposition 2.1.
-Then, for every divergence–free velocity field
+Let $T>0,\,s\in(0,\tfrac{1}{2})$ and $p>4$, and choose $\gamma,\lambda$ as in (2.8) and (2.11).
+Let $\overline{\delta}_{\theta}(s,p)$ and $\overline{\delta}_{u}(s,p)$ be given by (2.12).
+Then for every
+$0<\delta_{\theta}<\overline{\delta}_{\theta}(s,p),\qquad 0<\delta_{u}<\overline{\delta}_{u}(s,p),$
+the following holds. Set
+$\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2},\qquad\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4},$
+and let $Z_{t}^{\varepsilon}=-\sqrt{\varepsilon}\int_{0}^{t}\Delta e^{(t-r)\Delta}D\,dW_{r}$ be the continuous version in $H^{-2\alpha_{\theta}}(\mathcal{D})$ from Proposition 2.1.
+Then, for every divergence-free velocity field
 $u\in L^{p}\bigl(0,T;\prescript{}{0}{H}^{\frac{3}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\bigr)\cap L^{\infty}\bigl(0,T;\prescript{}{0}{H}^{\frac{1}{2}-\delta_{u}}_{\sigma}(\mathcal{D})\bigr)$
@@
 with trajectories in
-$\zeta^{\varepsilon}\in C(0,T;W^{s,6/5}(\mathcal{D}))\quad\mathbb{P}-\text{a.s.}$
+$\zeta^{\varepsilon}\in C(0,T;W^{s,6/5}(\mathcal{D}))\qquad\mathbb{P}\text{-a.s.}$
 A straightforward but important consequence of the proof of the previous result is the following estimate.
@@
 We now turn to the analysis of the 3D Navier-Stokes equations (2.3), assuming that $f^{\theta}$ is a prescribed forcing term.
-Global well-posedness results for the 3D Navier–Stokes equations are well known under smallness assumptions on the initial datum and the forcing; see for instance [8, Chapter 9].
+Global well-posedness results for the 3D Navier–Stokes equations are well known under smallness assumptions on the initial datum and the forcing; see for instance [9, Chapter 9].
 In the classical energy setting one typically assumes $f^{\theta}\in L^{2}(0,T;L^{2}(\mathcal{D}))$ and $u_{0}\in H^{1}(\mathcal{D})$.
@@
 Definition 2.7.
-The triple $(u^{\varepsilon},\theta^{\varepsilon},\tau^{\varepsilon})$ is a solution of (1.1) on $[0,T]$ if $\tau^{\varepsilon}\colon\Omega\to[0,T]$ is a stopping time and $(u^{\varepsilon},\theta^{\varepsilon})$ is a stochastic process with trajectories in
+The triple $(u^{\varepsilon},\theta^{\varepsilon},\tau^{\varepsilon})$ is a solution of (1.1) on $[0,T]$ if $\tau^{\varepsilon}\colon\Omega\to[0,T]$ is a stopping time and $(u^{\varepsilon},\theta^{\varepsilon})$ is a stochastic process with trajectories $\mathbb{P}$-a.s.
+in
 $\mathbb{E}^{\delta_{u}}_{\tau^{\varepsilon},p}\times C\bigl(0,\tau^{\varepsilon};H^{-\tfrac{1}{2}-\delta_{u}}(\mathcal{D})\bigr),$
@@
 Theorem 2.8.
-Let $T>0$ and $s\in[0,\tfrac{1}{2})$. Fix $p>4.$ Choose a parameter $\gamma>\frac{1}{4}$ sufficiently close to $\frac{1}{4}$ such that
-$a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}<1,\qquad\lambda_{\max}:=p(1-a(s,\gamma))>\lambda_{\min}:=\frac{1}{2}-s.$
-(Such a $\gamma$ always exists since $p>4$ implies p(1a(s,1/4))>1/2sp(1-a(s,1/4))>1/2-s). Choose
-$\lambda\in(\lambda_{\min},\,\lambda_{\max})\cap(0,1),$
-and define the thresholds
-$\overline{\delta}_{\theta}:=\min\{\lambda,\ 2\gamma-\tfrac{1}{2}\},\quad\overline{\delta}_{u}:=\min\{\lambda_{\max}-\lambda,1-\lambda\}.$
+Let $T>0$, $s\in(0,\tfrac{1}{2})$ and $p>4$, and choose $\gamma,\lambda$ as in (2.8) and (2.11).
+Let $\overline{\delta}_{\theta}$ and $\overline{\delta}_{u}$ be given by (2.12).
 Let parameters $\delta_{\theta},\delta_{u}$ satisfy
-$0<\delta_{\theta}\leq\overline{\delta}_{\theta},\qquad 0<\delta_{u}<\overline{\delta}_{u},$
+$0<\delta_{\theta}<\overline{\delta}_{\theta},\qquad 0<\delta_{u}<\overline{\delta}_{u},$
 subject to the compatibility conditions
-$\delta_{u}\geq\max\left\{\delta_{\theta},\ \tfrac{1}{2}-s\right\},\quad\delta_{u}<1-\frac{2}{p}.$
-Set
-$\alpha_{\theta}:=\tfrac{1}{4}+\tfrac{\delta_{\theta}}{2},\qquad\beta_{\theta}:=\tfrac{1}{4}-\tfrac{\delta_{\theta}}{4},$
+(2.15) $\delta_{u}\geq\max\left\{\delta_{\theta},\frac{1}{2}-s\right\},\qquad\delta_{u}<1-\frac{2}{p},$
+which define a non-empty set for the given range of $s$ and $p$. Set
+$\alpha_{\theta}:=\frac{1}{4}+\frac{\delta_{\theta}}{2},\qquad\beta_{\theta}:=\frac{1}{4}-\frac{\delta_{\theta}}{4},$
 and assume
@@
 $\mathbb{P}(\tau^{\varepsilon}=T)\geq 1-\frac{64\,\widetilde{M}^{2}\,\varepsilon}{\eta^{2}}\,C(\delta_{\theta},T)\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}.$
-Remark 2.9 (Consistency of parameters).
-Let us briefly verify that the set of admissible parameters is non-empty under the condition $p>4$.
-  1. (1)
-    Existence of $\lambda$.
-As $\gamma\downarrow\frac{1}{4}$, we have $a(s,\gamma)\to\frac{3}{4}+\frac{s}{2}$.
-Consequently, $\lambda_{\max}\to p(\frac{1}{4}-\frac{s}{2})$.
-The condition $\lambda_{\max}>\lambda_{\min}=\frac{1}{2}-s$ is equivalent to $p>\frac{4(1/2-s)}{1-2s}=2$.
-Since $p>4$, the interval for $\lambda$ is non-empty for $\gamma$ sufficiently close to $1/4$.
-  2. (2)
-    Existence of $\delta_{u}$.
-We need to ensure that the lower bound for $\delta_{u}$ is strictly smaller than the upper bound.
-Taking $\delta_{\theta}$ sufficiently small, the lower bound is $\frac{1}{2}-s$.
-On the other hand, the upper bound is $\min\{\lambda_{\max}-\lambda,1-\lambda,1-\frac{2}{p}\}$.
-We thus need to check that there exists a $\lambda$ such that
-    $\frac{1}{2}-s<\min\left\{1-\frac{2}{p},\ \lambda_{\max}-\lambda,\ 1-\lambda\right\}.$
-    • First, the condition $\frac{1}{2}-s<1-\frac{2}{p}$ is equivalent to $p>\frac{4}{1+2s}$.
-Since $s\in[0,1/2)$, the term $\frac{4}{1+2s}$ is at most $4$.
-Thus, the hypothesis $p>4$ guarantees this condition holds.
-    • Second, we check that $\frac{1}{2}-s<\lambda_{\max}-\lambda$.
-This is equivalent to finding $\lambda$ such that $\lambda<\lambda_{\max}-(\frac{1}{2}-s)$.
-Since we can choose $\lambda$ arbitrarily close to its lower bound $\lambda_{\min}=\frac{1}{2}-s$, such a $\lambda$ exists if $\lambda_{\min}<\lambda_{\max}-\lambda_{\min}$, i.e., $2\lambda_{\min}<\lambda_{\max}$.
-Since $\lambda_{\max}=p\bigl(1-a(s,\gamma)\bigr)$ and $a(s,\gamma)\to\frac{3}{4}+\frac{s}{2}$ as $\gamma\downarrow\frac{1}{4}$, we have $\lim\limits_{\gamma\downarrow\frac{1}{4}}\lambda_{\max}=\frac{p}{4}(1-2s);$ note that $\frac{p}{4}(1-2s)>2\lambda_{\min}=1-2s$, thanks to the hypothesis $p>4$.
-Thus the original condition is verified.
-    • Finally, we check that $\frac{1}{2}-s<1-\lambda$.
-This is equivalent to finding $\lambda$ such that $\lambda<\frac{1}{2}+s$.
-Since $\lambda$ can be chosen close to $\lambda_{\min}=\frac{1}{2}-s$, i.e.
-$\lambda=\lambda_{\min}+\varepsilon_{\lambda},$ with $\varepsilon_{\lambda}>0$.
-Thus the condition can be rewritten as $\frac{1}{2}-s<\frac{1}{2}+s+\varepsilon_{\lambda}$, which holds for any $s\geq 0$.
 2.4. Overview
@@
 We start this section by proving the first part of Proposition 2.1, in particular the well-posedness and the regularity for the mild solution $Z_{t}^{\varepsilon}$ of the linear problem (2.4) subject to Dirichlet noise boundary conditions.
-Note that this result is well-known in the literature, see for instance [9, Proposition 3.1].
+Note that this result is well-known in the literature, see for instance [10, Proposition 3.1].
 Proof of Proposition 2.1 (i)-(ii).
@@
 (3.2) $\mathbb{E}\left[\norm{\xi_{t_{2}}-\xi_{t_{1}}}_{2}^{2}\right]\leq M(t_{2}-t_{1})^{\eta},\quad\forall\,0\leq t_{1}\leq t_{2},$
-then the thesis follows from the Kolmogorov test for Gaussian processes, see [11, Proposition 3.16]. Consider
+then the thesis follows from the Kolmogorov test for Gaussian processes, see [12, Proposition 3.16]. Consider
 $\quad\xi_{t}=\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}}e^{(t-r)\Delta}DdW_{r}=\int_{0}^{t}(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{(t-r)\Delta}(-\Delta)^{\beta_{\theta}}DdW_{r},$
@@
 (3.4) $\begin{split}\mathbb{E}\left[\|I_{1}\|_{2}^{2}\right]\leq C(t_{2}-t_{1})^{\frac{\delta_{\theta}}{2}}\sum_{k}\lambda_{k}^{2}\norm{(-\Delta)^{\beta_{\theta}}De_{k}}_{2}^{2}.\end{split}$
-For the estimate of the second integral $I_{2}$, we use the fact that for any $\gamma\in(0,\alpha_{\theta}+\beta_{\theta})$ there exists $C>0$ such that for any $0\leq r\leq t_{1}$, it holds
-(3.5) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2}(\mathcal{D}))}\leq C(t_{2}-t_{1})^{\gamma}(t_{1}-r)^{-(1-\alpha_{\theta}-\beta_{\theta}+\gamma)},$
-where $\mathcal{L}(L^{2})$ is the space of bounded, linear operators from $L^{2}(\mathcal{D})$ to itself. Indeed, by the semigroup property
-$e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta}=e^{(t_{1}-r)\Delta}(e^{(t_{2}-t_{1})\Delta}-I).$
-Further, by the commuting property of the fractional powers with the heat semigroup
-$(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})=(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I).$
-Considering the norms on the previous identity, it yields
-(3.6) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2})}\leq\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|_{\mathcal{L}(L^{2})}\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}.$
-The first term on the right-hand-side can be estimated by the analyticity of the semigroup as
-(3.7) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|_{\mathcal{L}(L^{2})}\leq C(t_{1}-r)^{-1+\alpha_{\theta}+\beta_{\theta}-\gamma}.$
-For the second term on the right-hand side, again thanks to the analyticity of the semigroup, it holds
-$(e^{(t_{2}-t_{1})\Delta}-I)g=\Delta\int_{0}^{t_{2}-t_{1}}e^{\sigma\Delta}g\,d\sigma$
-for any $g\in L^{2}(\mathcal{D})$, see [22, Proposition 2.1.4].
-Applying $(-\Delta)^{-\gamma}$, using $1-\gamma<1$, the fact that $(-\Delta)^{1-\gamma}$ is a closed operator and $\sigma\mapsto(-\Delta)^{1-\gamma}e^{\sigma\Delta}$ is integrable for any $g\in L^{2}(\mathcal{D})$, we obtain, thanks to [21, Proposition A.5],
-$(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)g=-\int_{0}^{t_{2}-t_{1}}(-\Delta)^{1-\gamma}e^{\sigma\Delta}g\,d\sigma.$
-Again by the estimate for the fractional powers, we have
-$\begin{split}\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)g\|_{2}&\leq\int_{0}^{t_{2}-t_{1}}\|(-\Delta)^{1-\gamma}e^{\sigma\Delta}g\|_{2}\,d\sigma\\ &\leq C\int_{0}^{t_{2}-t_{1}}\sigma^{-1+\gamma}\|g\|_{2}\,d\sigma\\ &=\frac{C}{\gamma}(t_{2}-t_{1})^{\gamma}\|g\|_{2}.\end{split}$
-Since the previous estimate holds for any $g\in L^{2}(\mathcal{D})$, we have checked
-(3.8) $\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}\leq\frac{C}{\gamma}(t_{2}-t_{1})^{\gamma}$
-for any $\gamma\in(0,\alpha_{\theta}+\beta_{\theta}).$ Substituting (3.7) and (3.8) into (3.6) gives (3.5).
-Thus, to estimate $I_{2}$, using the independence of $(\beta_{k})_{k}$, Itô-isometry and (3.5), we obtain, for any $\gamma\in(0,\alpha_{\theta}+\beta_{\theta})$,
-(3.9) $\begin{split}\mathbb{E}\left[\|I_{2}\|_{2}^{2}\right]&\leq\sum_{k}\lambda_{k}^{2}\int_{0}^{t_{1}}\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}\left[e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta}\right](-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,dr\\ &\leq C\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,(t_{2}-t_{1})^{2\gamma}\int_{0}^{t_{1}}(t_{1}-r)^{-2(1-\alpha_{\theta}-\beta_{\theta}+\gamma)}\,dr.\end{split}$
-Note that the integral on the right-hand side is finite if $1-\alpha_{\theta}-\beta_{\theta}+\gamma<\frac{1}{2},$ i.e.
-$\gamma<\alpha_{\theta}+\beta_{\theta}-\frac{1}{2}=\frac{\delta_{\theta}}{4}$.
-Thus, choosing $\gamma\in(0,\frac{\delta_{\theta}}{4})$, we arrive to
-(3.10) $\mathbb{E}\left[\|I_{2}\|_{2}^{2}\right]\leq C\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}(t_{2}-t_{1})^{2\gamma},$
+For the estimate of the second integral $I_{2}$, fix $0\leq t_{1}\leq t_{2}$. By the semigroup property,
+$e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta}=e^{(t_{1}-r)\Delta}(e^{(t_{2}-t_{1})\Delta}-I),\qquad 0\leq r\leq t_{1}.$
+Let $\gamma\in(0,1]$ (to be chosen later). Using the commutation of $(-\Delta)^{\sigma}$ with $e^{t\Delta}$ and inserting $(-\Delta)^{\gamma}(-\Delta)^{-\gamma}$, we obtain
+(3.5) $\begin{split}\norm{(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})}_{\mathcal{L}(L^{2})}&=\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}e^{(t_{1}-r)\Delta}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}\\ &=\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}\\ &\leq\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|\,\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}.\end{split}$
+For the first factor we use [25, Theorem 6.13(c)] and we get
+(3.6) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}+\gamma}e^{(t_{1}-r)\Delta}\|_{\mathcal{L}(L^{2})}\leq C(t_{1}-r)^{-(1-\alpha_{\theta}-\beta_{\theta}+\gamma)},$
+for a constant $C$ independent of $t_{1}$.
+For the second factor, let $g\in L^{2}(\mathcal{D})$ and set $x:=(-\Delta)^{-\gamma}g\in D((-\Delta)^{\gamma})$.
+Since $(-\Delta)^{-\gamma}$ commutes with the heat semigroup $e^{(t_{2}-t_{1})\Delta}$, we have
+$(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)g=(e^{(t_{2}-t_{1})\Delta}-I)(-\Delta)^{-\gamma}g=(e^{(t_{2}-t_{1})\Delta}-I)x.$
+Applying [25, Theorem 6.13(d)] with exponent $\gamma$ yields
+$\|(e^{(t_{2}-t_{1})\Delta}-I)x\|_{2}\leq C(t_{2}-t_{1})^{\gamma}\|(-\Delta)^{\gamma}x\|_{2}=C(t_{2}-t_{1})^{\gamma}\|g\|_{2},$
+for a new constant $C$ independent of $t_{1},t_{2}$. Therefore
+(3.7) $\|(-\Delta)^{-\gamma}(e^{(t_{2}-t_{1})\Delta}-I)\|_{\mathcal{L}(L^{2})}\leq C(t_{2}-t_{1})^{\gamma}.$
+Combining (3.5)-(3.7) we obtain, for all $0\leq r\leq t_{1}$,
+(3.8) $\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2})}\leq C(t_{2}-t_{1})^{\gamma}(t_{1}-r)^{-(1-\alpha_{\theta}-\beta_{\theta}+\gamma)}.$
+Using independence of $(\beta_{k})_{k}$, Itô isometry, and (3.8), we get
+$\mathbb{E}\|I_{2}\|_{2}^{2}$ $=\sum_{k}\lambda_{k}^{2}\int_{0}^{t_{1}}\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,dr$
+$\leq\sum_{k}\lambda_{k}^{2}\int_{0}^{t_{1}}\|(-\Delta)^{1-\alpha_{\theta}-\beta_{\theta}}(e^{(t_{2}-r)\Delta}-e^{(t_{1}-r)\Delta})\|_{\mathcal{L}(L^{2})}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\,dr$
+$\leq C(t_{2}-t_{1})^{2\gamma}\left(\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2}\right)\int_{0}^{t_{1}}(t_{1}-r)^{-2(1-\alpha_{\theta}-\beta_{\theta}+\gamma)}\,dr.$
+The time-integral is finite provided $2(1-\alpha_{\theta}-\beta_{\theta}+\gamma)<1$, i.e.
+$\gamma<\alpha_{\theta}+\beta_{\theta}-\frac{1}{2}.$
+Since $\alpha_{\theta}+\beta_{\theta}=\frac{1}{2}+\frac{\delta_{\theta}}{4}$, we may choose any $\gamma\in(0,\tfrac{\delta_{\theta}}{4})$, and then
+(3.9) $\mathbb{E}\|I_{2}\|_{2}^{2}\leq C\,(t_{2}-t_{1})^{2\gamma}\sum_{k}\lambda_{k}^{2}\|(-\Delta)^{\beta_{\theta}}De_{k}\|_{2}^{2},$
 where $C>0$ depends on $T,\alpha_{\theta},\beta_{\theta},\gamma$ but not on $t_{1},t_{2}$.
@@
 $Y_{\rho}(t):=\int_{0}^{t}(t-r)^{-\rho}e^{(t-r)\Delta}\Phi dW_{r},\quad t\in[0,T].$
-By the factorisation method, see [11, Section 5.3.1], since $\int_{0}^{T}t^{-2\rho}\|e^{t\Delta}\Phi\|_{L^{0}_{2}}^{2}dt<\infty$, it holds
+By the factorisation method, see [12, Section 5.3.1], since $\int_{0}^{T}t^{-2\rho}\|e^{t\Delta}\Phi\|_{L^{0}_{2}}^{2}dt<\infty$, it holds
 $W^{\Phi}_{\Delta}(t)=\frac{\sin(\pi\rho)}{\pi}\int_{0}^{t}(t-r)^{\rho-1}e^{(t-r)\Delta}Y_{\rho}(r)\,dr.$
@@
 $\overline{\delta}_{\theta}(\lambda,\gamma):=\min\{\lambda,2\gamma-\frac{1}{2}\}.$
-For any $0<\delta_{\theta}\leq\overline{\delta}_{\theta}$, any velocity field
+For any $0<\delta_{\theta}<\overline{\delta}_{\theta}$, any velocity field
 $u\in L^{p}(0,T;H^{\frac{3}{2}-\delta_{u}}(\mathcal{D}))\cap L^{\infty}(0,T;H^{\frac{1}{2}-\delta_{u}}(\mathcal{D})),\qquad p>1,\,\delta_{u}\in(0,1-\lambda],$
@@
 To simplify the notation, we set $\varepsilon=1$ and drop the dependence on $\varepsilon$ of $\xi^{\varepsilon}$.
-We start by recalling that from [6, Theorem 7.4], given $s\geq 0$, the pointwise multiplication
+We start by recalling that, by [6, Theorem A.1] (which we can apply after extending functions on $\mathcal{D}$ to $\mathbb{R}^{3}$ via a bounded extension operator), the pointwise multiplication
 $W^{s_{1},p_{1}}\left(\mathcal{D}\right)\times W^{s_{2},p_{2}}\left(\mathcal{D}\right)$ $\hookrightarrow W^{s,p}\left(\mathcal{D}\right)$
@@
   3. (iii)
-    $s_{1}+s_{2}>s+3\left(\frac{1}{p_{1}}+\frac{1}{p_{2}}-\frac{1}{p}\right).$
+    $s_{1}+s_{2}>s+3\left(\frac{1}{p_{1}}+\frac{1}{p_{2}}-\frac{1}{p}\right)$;
+  4. (iv)
+    if $s_{i}=s\notin\mathbb{N}$ for some $i\in\{1,2\}$, then one must additionally assume $p_{i}\leq p$.
 Step 1. As a first step of the proof, we show that
@@
 By the norm equivalence between $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})$ and $H^{\frac{1}{2}+\lambda}(\mathcal{D})$, and by Corollary 3.1, we have $u_{t}\in W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)$ a.e., and since $\varphi\in L^{6}\left(\mathcal{D}\right)$, we deduce $\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\in W^{2\gamma,6}\left(\mathcal{D}\right)$.
-Hence, (3.15) follows from the pointwise multiplication embedding
+Thus (3.14) follows from the pointwise multiplication embedding
 $W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)\times W^{2\gamma,6}\left(\mathcal{D}\right)$ $\hookrightarrow W^{\frac{1}{2}+\delta_{\theta},2}\left(\mathcal{D}\right)$
@@
     $\delta_{\theta}<\lambda+2\gamma-\frac{1}{2}.$
+Since in general $s=\frac{1}{2}+\delta_{\theta}\notin\mathbb{N}$ and here $p_{2}=6>p=2$, condition (iv) in [6, Theorem A.1] forbids the borderline case $s_{2}=s$. Hence we impose
+$s_{2}>s\quad\Longleftrightarrow\quad 2\gamma>\frac{1}{2}+\delta_{\theta}\quad\Longleftrightarrow\quad\delta_{\theta}<2\gamma-\frac{1}{2}.$
+In contrast, the endpoint $\delta_{\theta}=\lambda$ is admissible since then $s_{1}=s$ but $p_{1}=p=2$.
 Therefore, we can choose
@@
 Corollary 3.3.
-Let $s\in[0,\frac{1}{2})$, $p>1$ and $\gamma>\frac{1}{4}$ be given. Assume that
+Let $s\in(0,\frac{1}{2})$, $p>1$ and $\gamma>\frac{1}{4}$ be given. Assume that
 (3.16) $a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}<1.$
@@
 (3.17) $\lambda<p(1-a(s,\gamma)).$
-Fix $0<\delta_{\theta}\leq\overline{\delta}_{\theta}=\min\{\lambda,2\gamma-\frac{1}{2}\},$ and consider $\alpha_{\theta}:=\frac{1}{4}+\frac{\delta_{\theta}}{2}.$ Define
+Fix $0<\delta_{\theta}<\overline{\delta}_{\theta}=\min\{\lambda,2\gamma-\frac{1}{2}\},$ and consider $\alpha_{\theta}:=\frac{1}{4}+\frac{\delta_{\theta}}{2}.$ Define
 $\overline{\delta}_{u}(p,s,\gamma,\lambda):=\min\{p(1-a(s,\gamma))-\lambda,1-\lambda\}>0.$
@@
 where $C=C(s,p,\gamma,\lambda,\delta_{\theta},T)>0$.
-Lastly, given $s\in[0,\frac{1}{2}),p>1$ as before, there always exists $\gamma=\gamma(s)>\frac{1}{4}$ and $\lambda=\lambda(s,p)\in(0,1)$ such that (3.17) and (3.18) hold.
+Lastly, given $s\in(0,\frac{1}{2}),p>1$ as before, there always exists $\gamma=\gamma(s)>\frac{1}{4}$ and $\lambda=\lambda(s,p)\in(0,1)$ such that (3.16) and (3.17) hold.
 Proof.
@@
 We deduce that $\|\rho_{t}-\rho_{r}\|_{L^{\frac{6}{5}}(\mathcal{D})}\to 0$ for every $r\in(0,T]$, that is $\rho\in C(0,T;L^{\frac{6}{5}}(\mathcal{D}))$.
-Consequently, $\chi_{t}=(-\Delta)^{-\frac{s}{2}}\rho_{t}$ belongs to $C(0,T;W^{s,\frac{6}{5}}(\mathcal{D})\bigr)$.
+Consequently, $\chi_{t}=(-\Delta)^{-\frac{s}{2}}\rho_{t}$ belongs to $C(0,T;W^{s,\frac{6}{5}}(\mathcal{D}))$.
 Step 4.
 Existence of $\gamma,\lambda$, given $s,p$.
-Assume that, as before, $s\in[0,\frac{1}{2})$ and $p>1$ are given.
+Assume that, as before, $s\in(0,\frac{1}{2})$ and $p>1$ are given.
 We now check that we can always find $\gamma=\gamma(s)>\frac{1}{4}$ and $\lambda=\lambda(s,p)$ such that (3.16) and (3.17) are satisfied, and thus this corollary is non-trivial.
@@
 Remark 3.4.
-From the proof of Corollary 3.3, it follows that if $s\in[0,\frac{1}{2}),p>1$, $\gamma>\frac{1}{4}$ and $\lambda\in(0,1)$ satisfy (3.17) and (3.18), then
+From the proof of Corollary 3.3, it follows that if $s\in(0,\frac{1}{2}),p>1$, $\gamma>\frac{1}{4}$ and $\lambda\in(0,1)$ satisfy (3.16) and (3.17), then
 $\bigl(\gamma+\tfrac{s}{2}+\tfrac{1}{2}\bigr)\,\frac{1}{1-\frac{\lambda}{p}-\frac{\delta_{u}}{p}}<1$
@@
 Proof of Theorem 2.2.
-For simplicity we set $\varepsilon=1$ and drop the superscript, writing $\zeta$ and $Z$ instead of $\zeta^{\varepsilon}$ and $Z^{\varepsilon}$.
-We also use the notation of Corollary 3.3 and Proposition 2.1.
-Let $s\in[0,\frac{1}{2})$ and $p>2$ be given.
-Step $1$. Choices of $\gamma,q,q^{\prime},\lambda,\overline{\delta}_{\theta},\overline{\delta}_{u}.$
-Choice of $\gamma=\gamma(s,p)>\frac{1}{4}.$ We choose $\gamma=\frac{1}{4}+\varepsilon_{\gamma}$, with $\varepsilon_{\gamma}>0$ sufficiently small such that
-(3.22) $\left\{\begin{aligned} a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}&<1,\\ \lambda_{\max}(s,p,\gamma):=p(1-a(s,\gamma))&>\frac{1}{2}-s.\end{aligned}\right$.
-The condition (3.22)1 is satisfied, up to choosing $\varepsilon_{\gamma}$ sufficiently small, since
-$\lim\limits_{\gamma\downarrow\frac{1}{4}}a(s,\gamma)=\frac{3}{4}+\frac{s}{2}<1,$
-where in the last inequality we use $s<\frac{1}{2}$.
-For the condition (3.22)2, note that if $\gamma\in(\frac{1}{4},\frac{1-s}{2})$, then $0<\frac{1}{2}-\gamma-\frac{s}{2}=1-a(s,\gamma)$.
-Further, since $p>2$ by hypothesis, we have
-$\lim\limits_{\gamma\downarrow\frac{1}{4}}\lambda_{\max}(s,p,\gamma)=\frac{p}{2}(\frac{1}{2}-s)>\frac{1}{2}-s.$
-Choice of $q=q(s)$ and $q^{\prime}=q^{\prime}(s)$. Let $q=q(s):=\frac{6}{5-2s}.$ With this choice, the Sobolev embedding
-(3.23) $W^{s,\frac{6}{5}}(\mathcal{D})\hookrightarrow L^{q(s)}(\mathcal{D})$
-holds, a condition that we will later need in the proof. Further, denoting by $q^{\prime}$ the conjugate exponent of $q$, it holds
-(3.24) $1-\frac{3}{q^{\prime}}<\lambda_{\max}(s,\gamma,p).$
-Indeed
-$1-\frac{3}{q^{\prime}(s)}=-2+\frac{3}{q(s)}=\frac{1}{2}-s,$
-and (3.24) then follows from (3.22)2.
-Choice of $\lambda=\lambda(s).$ Define
-$\lambda_{\min}(s):=\frac{1}{2}-s=1-\frac{3}{q^{\prime}(s)}.$
-and note that $\lambda_{\min}\in(0,1)$ since $s\in[0,\frac{1}{2})$. Further, $\lambda_{\min}<\lambda_{\max}$. Thus, we can choose $\lambda$ sufficiently close to $\lambda_{\min}$ such that
-$\lambda=\lambda(s)\in(\lambda_{\min},\lambda_{\max})\cap(0,1).$
-With this choice of $\lambda$, we claim that the pointwise multiplication embedding
-(3.25) $W^{\frac{1}{2}+\lambda(s),2}(\mathcal{D})\times W^{2\gamma(s),6}(\mathcal{D})\hookrightarrow L^{q^{\prime}(s)}(\mathcal{D})$
-holds, another condition that we will need later in the proof. Indeed, thanks to [6, Theorem 7.4], it is sufficient to check that
-  1. (i)
-    $\frac{1}{2}+\lambda\geq 0$, $2\gamma\geq 0$;
-  2. (ii)
-    $\frac{1}{2}+\lambda-\frac{3}{2}\geq-\frac{3}{q^{\prime}}$, $2\gamma-\frac{1}{2}\geq-\frac{3}{q^{\prime}}$;
-  3. (iii)
-    $\frac{1}{2}+\lambda+2\gamma-3(\frac{1}{2}+\frac{1}{6})>-\frac{3}{q^{\prime}}.$
-The previous conditions are equivalent to
-  1. (a)
-    $\lambda\geq-\frac{1}{2},$ $\gamma\geq 0$;
-  2. (b)
-    $\lambda\geq 1-\frac{3}{q^{\prime}},$ $2\gamma\geq\frac{1}{2}-\frac{3}{q^{\prime}};$
-  3. (c)
-    $\lambda+2\gamma\geq\frac{3}{2}-\frac{3}{q^{\prime}}$.
-The conditions in (a) and (b) follow from the choice of $\lambda,\gamma$, in particular thanks to $\lambda>\lambda_{\min}>0$ and $\gamma>\frac{1}{4}.$ Lastly, since $\lambda+2\gamma>\lambda+\frac{1}{2}$, then (c) is satisfied if $\lambda+\frac{1}{2}>\frac{3}{2}-\frac{3}{q^{\prime}}$, which is true thanks to the choice $\lambda>\lambda_{\min}.$
-Choice of $\overline{\delta}_{\theta}$ and $\overline{\delta}_{u}$. With $\gamma$ and $\lambda$ fixed as above, we define the thresholds
-$\overline{\delta}_{\theta}(s,p):=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\},\quad\overline{\delta}_{u}(s,p):=\min\left\{\lambda_{\max}(s,p,\gamma)-\lambda,1-\lambda\right\},$
-and note that $\overline{\delta}_{\theta}$ and $\overline{\delta}_{u}$ are the same as in Corollary 3.3.
-Further, both are positive by the choices made above.
-We assume $0<\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $0<\delta_{u}<\overline{\delta}_{u}$.
-Since $\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $\delta_{u}<\overline{\delta}_{u}$, the assumptions of Corollary 3.3 are satisfied.
-Furthermore, from Remark 3.4, it follows that
-(3.26) $(\gamma+\frac{s}{2}+\frac{1}{2})\frac{1}{1-\frac{\lambda}{p}-\frac{\delta_{u}}{p}}<1.$
+For simplicity we set $\varepsilon=1$ and drop the superscript, writing $\zeta$ and $Z$ instead of $\zeta^{\varepsilon}$ and $Z^{\varepsilon}$. Let $s\in(0,\frac{1}{2})$ and $p>4$ be given.
+Step 1.
+Setup of parameters and auxiliary embeddings.
+We work with the parameters $\gamma,\lambda,\overline{\delta}_{\theta},\overline{\delta}_{u}$ fixed in the statement of the theorem.
+In particular,
+$a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}<1,\qquad\lambda_{\max}:=p(1-a(s,\gamma)),$
+and, by (2.9), we have $\lambda_{\max}>\frac{1}{2}-s$. Moreover, $\lambda_{\max}-(1-2s)>0$ by the choice of $\varepsilon_{\gamma}$ in (2.8) as shown in (2.10). We have set
+$\lambda:=(\frac{1}{2}-s)+\varepsilon_{\lambda},\qquad 0<\varepsilon_{\lambda}<\min\{2s,\lambda_{\max}-(1-2s)\}.$
+Note that
+$\lambda<(\tfrac{1}{2}-s)+2s=\tfrac{1}{2}+s<1,\qquad\lambda<(\tfrac{1}{2}-s)+\lambda_{\max}-(1-2s)=\lambda_{\max}-(\tfrac{1}{2}-s)<\lambda_{\max}.$
+Hence $\lambda\in(\tfrac{1}{2}-s,\min\{1,\lambda_{\max}\})$. The thresholds are
+$\overline{\delta}_{\theta}=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\},\qquad\overline{\delta}_{u}=\min\left\{\lambda_{\max}-\lambda,1-\lambda\right\}.$
+Finally, fix $0<\delta_{\theta}<\overline{\delta}_{\theta}$ and $0<\delta_{u}<\overline{\delta}_{u}$.
+Choice of $q,q^{\prime}$. Set
+$q:=\frac{6}{5-2s},\qquad\frac{1}{q}+\frac{1}{q^{\prime}}=1.$
+Then the Sobolev embedding
+(3.21) $W^{s,\frac{6}{5}}(\mathcal{D})\hookrightarrow L^{q}(\mathcal{D})$
+holds, and a direct computation gives
+(3.22) $1-\frac{3}{q^{\prime}}=\frac{1}{2}-s.$
+Product embedding. Since $\gamma>\frac{1}{4}$, the Sobolev embedding yields
+(3.23) $W^{2\gamma,6}(\mathcal{D})\hookrightarrow L^{\infty}(\mathcal{D}).$
+Moreover, again by the Sobolev embedding,
+(3.24) $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\hookrightarrow L^{r_{\lambda}}(\mathcal{D}),\qquad r_{\lambda}:=\frac{3}{1-\lambda}.$
+Recalling that $q^{\prime}=\tfrac{6}{1+2s},$ it can be checked that $r_{\lambda}\geq q^{\prime}$ is equivalent to $\lambda\geq\tfrac{1}{2}-s$.
+Hence, by (3.24) and the embedding $L^{r_{\lambda}}(\mathcal{D})\hookrightarrow L^{q^{\prime}}(\mathcal{D})$ on bounded $\mathcal{D}$,
+(3.25) $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\hookrightarrow L^{q^{\prime}}(\mathcal{D}).$
+Combining (3.23), (3.25) with Hölder’s inequality yields the continuous multiplication map
+(3.26) $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\times W^{2\gamma,6}(\mathcal{D})\hookrightarrow L^{q^{\prime}}(\mathcal{D}).$
+In particular, for every $\varphi\in L^{6}(\mathcal{D})$,
+(3.27) $\bigl\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\bigr\|_{L^{q^{\prime}}(\mathcal{D})}\leq C\,\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\,\|\varphi\|_{L^{6}(\mathcal{D})}.$
+Time integrability condition. Since $\delta_{u}<\overline{\delta}_{u}\leq 1-\lambda$, we have $\lambda\leq 1-\delta_{u}$ and therefore Corollary 3.1 implies
+(3.28) $u\in L^{\frac{p}{\lambda+\delta_{u}}}\bigl(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\bigr).$
+Moreover, $\delta_{u}<\overline{\delta}_{u}\leq\lambda_{\max}-\lambda$ gives $\lambda+\delta_{u}<\lambda_{\max}=p(1-a(s,\gamma))$, and hence, by Remark 3.4,
+(3.29) $\Bigl(\gamma+\frac{s}{2}+\frac{1}{2}\Bigr)\frac{1}{1-\frac{\lambda}{p}-\frac{\delta_{u}}{p}}<1.$
 Step 2. Fixed point argument. Observe that, using the notation introduced in Corollary 3.3, we can write $\zeta_{t}$ in the form
@@
 Step 3. We want to show that, for any fixed $0<\delta_{u}<\overline{\delta}_{u}(s,p)$ and any $u$ as in the statement of the theorem, it holds
-$(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T_{1};L^{\frac{6}{5}}(\mathcal{D})).$
-As in the proof of Lemma 3.2, given $\varphi\in L^{6}(\mathcal{D})$ we have
-$\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle$ $=\langle\mathrm{div}(u_{t}\zeta_{t}),(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle$
-$=-\langle u_{t}\zeta_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle$
-$=-\langle\zeta_{t},u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle.$
-We now consider the conjugate exponents $q=q(s),q^{\prime}=q^{\prime}(s)\in(1,+\infty)$ chosen in Step 1.
-We know that $u\in L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)$ thanks to the norm equivalence between $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})$ and $H^{\frac{1}{2}+\lambda}(\mathcal{D})$, and Corollary 3.1 (note that its hypotheses are satisfied since $\delta_{u}\in(0,1)$, $\lambda\in(0,1-\delta_{u})$).
-Since the pointwise multiplication embedding
-(3.28) $\begin{split}W^{\frac{1}{2}+\lambda(s),2}(\mathcal{D})\times W^{2\gamma(s),6}(\mathcal{D})&\hookrightarrow L^{q^{\prime}(s)}(\mathcal{D})\\ (u_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi)&\mapsto u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\end{split}$
-holds thanks to the condition verified after (3.25), we get that there exists $C=C(\lambda(s),\gamma(s))=C(s)>0$ such that
-$\left\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{L^{q^{\prime}}(\mathcal{D})}\leq C\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\left\|\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{W^{2\gamma,6}(\mathcal{D})},$
-and thus
-$\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}$ $=\sup_{\|\varphi\|_{6}\leq 1}\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle$
-$\leq C\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\sup_{\|\varphi\|_{6}\leq 1}\left\|\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{W^{2\gamma,6}(\mathcal{D})}$
-$\leq C\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\sup_{\|\varphi\|_{6}\leq 1}\left\|\varphi\right\|_{L^{6}(\mathcal{D})}$
-$=C\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}.$
-Further, since $\zeta\in X=C(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D}))\hookrightarrow C(0,T_{1};L^{q}(\mathcal{D}))$ thanks to (3.23), we can write
-(3.29) $\int_{0}^{T_{1}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dt\leq C\left(\sup_{t\in[0,T_{1}]}\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\right)^{\frac{p}{\lambda+\delta_{u}}}\int_{0}^{T_{1}}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dt<\infty.$
-The thesis of this step follows at this point since $u\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))$.
+(3.31) $(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\in L^{\frac{p}{\lambda+\delta_{u}}}\bigl(0,T_{1};L^{\frac{6}{5}}(\mathcal{D})\bigr).$
+Let $\varphi\in L^{6}(\mathcal{D})$. As in the proof of Lemma 3.2, using integration by parts and the self-adjointness of $\Delta$, we obtain
+$\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle$ $=\left\langle\mathrm{div}(u_{t}\zeta_{t}),(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle$
+$=-\left\langle u_{t}\zeta_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle=-\left\langle\zeta_{t},u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle.$
+Let $q,q^{\prime}$ be the conjugate exponents introduced in Step 1, namely $q=\frac{6}{5-2s}$ and $1/q+1/q^{\prime}=1$. By Hölder’s inequality,
+$\bigl|\left\langle\zeta_{t},u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle\bigr|\leq\|\zeta_{t}\|_{L^{q}(\mathcal{D})}\;\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\|_{L^{q^{\prime}}(\mathcal{D})}.$
+Using the product estimate (3.27) from Step 1, we infer
+$\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\|_{L^{q^{\prime}}(\mathcal{D})}\leq C\,\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\,\|\varphi\|_{L^{6}(\mathcal{D})},$
+and therefore
+$\bigl|\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle\bigr|\leq C\,\|\zeta_{t}\|_{L^{q}(\mathcal{D})}\;\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\;\|\varphi\|_{L^{6}(\mathcal{D})}.$
+Taking the supremum over $\|\varphi\|_{L^{6}}\leq 1$ yields
+(3.32) $\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}(u_{t}\zeta_{t})\|_{L^{\frac{6}{5}}(\mathcal{D})}\leq C\,\|\zeta_{t}\|_{L^{q}(\mathcal{D})}\;\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}.$
+Since $\zeta\in X=C(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D}))$, the Sobolev embedding (3.21) implies $\sup_{t\in[0,T_{1}]}\|\zeta_{t}\|_{L^{q}}<\infty$.
+Hence, raising (3.32) to the power $\frac{p}{\lambda+\delta_{u}}$ and integrating, we obtain
+(3.33) $\int_{0}^{T_{1}}\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}(u_{t}\zeta_{t})\|_{L^{\frac{6}{5}}}^{\frac{p}{\lambda+\delta_{u}}}\,dt\leq C\Bigl(\sup_{t\in[0,T_{1}]}\|\zeta_{t}\|_{L^{q}}\Bigr)^{\frac{p}{\lambda+\delta_{u}}}\int_{0}^{T_{1}}\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}}^{\frac{p}{\lambda+\delta_{u}}}\,dt.$
+The last integral is finite by (3.28), i.e. $u\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))$. This proves (3.31)
 Step 4. Repeating the same arguments in Corollary 3.3, we obtain $\Lambda\left(\zeta\right)_{t}\in C\left(0,T;L^{\frac{6}{5}}(\mathcal{D})\right)$, and thus $\Psi(\zeta)=\tilde{\zeta}\in X$.
 Step 5.
-We conclude the proof by showing that $\Psi$ is a contraction from $X$ to itself, up to choosing $T_{1}$ small enough.
-Consider $\zeta_{1},\zeta_{2}\in X$ and their difference $\zeta:=\zeta_{1}-\zeta_{2}$, then, from
+We conclude the proof by showing that $\Psi$ is a contraction in $X=C(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D}))$ for $T_{1}>0$ sufficiently small.
+Let $\zeta_{1},\zeta_{2}\in X$ and set $\zeta:=\zeta_{1}-\zeta_{2}$.
+From
 $\Psi(\zeta_{i})=e^{t\Delta}\theta_{0}-(-\Delta)^{-\frac{s}{2}}\Lambda(\zeta_{i})_{t}-\chi_{t},$
-we deduce
-$\Psi(\zeta_{1})-\Psi(\zeta_{2})=-(-\Delta)^{-\frac{s}{2}}\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)dr.$
-Considering the norm, we obtain
-(3.30) $\begin{split}\norm{\Psi(\zeta_{1})-\Psi(\zeta_{2})}_{X}&\leq\sup_{t\in[0,T_{1}]}\norm{\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)dr}_{L^{\frac{6}{5}}(\mathcal{D})}\\ &\leq\int_{0}^{T_{1}}\norm{(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(T_{1}-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)}_{L^{\frac{6}{5}}(\mathcal{D})}dr.\end{split}$
-Using the same argument applied in Step 2 of the proof of Corollary 3.3, we get
-(3.31) $\begin{split}\norm{\Psi(\zeta_{1})-\Psi(\zeta_{2})}_{X}&\leq\int_{0}^{T_{1}}\frac{c}{\left(T_{1}-r\right)^{\gamma+\frac{s}{2}+\frac{1}{2}}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}dr\\ &\leq\left(\int_{0}^{T_{1}}\left(\frac{c}{\left(T_{1}-r\right)^{\gamma+\frac{s}{2}+\frac{1}{2}}}\right)^{\frac{p}{p-\lambda-\delta_{u}}}dr\right)^{\frac{p-\lambda-\delta_{u}}{p}}\\ &\quad\cdot\left(\int_{0}^{T_{1}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\,\mathrm{div}(u_{r}\zeta_{r})\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dr\right)^{\frac{\lambda+\delta_{u}}{p}}.\end{split}$
-where $c=c(\gamma(s),s)=c(s)$. The first term is finite thanks to condition (3.26). The second term can be bounded as at the end of Step 3 in (3.29), to obtain
-$\begin{split}\left(\int_{0}^{T_{1}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dr\right)^{\frac{\lambda+\delta_{u}}{p}}\leq C\sup_{t\in[0,T_{1}]}\norm{\zeta}_{L^{q}(\mathcal{D})}\norm{u}_{L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)}.\end{split}$
-Here, $C=C(\lambda(s),\gamma(s))=C(s)$ is independent of the time variable. In conclusion, we have shown
-(3.32) $\norm{\Psi(\zeta_{1})-\Psi(\zeta_{2})}_{C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)}\leq K\,\norm{u}_{L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)}\,\norm{\zeta}_{C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)},$
-where we observe that $K$ is independent of the time variable.
-Thus, since $u\in L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)$, then $\Psi$ is a contraction from $X$ to itself if we choose $T_{1}>0$ sufficiently small.
-This gives the existence of a solution $\zeta\in C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)$.
-If $T_{1}<T$, we can iteratively repeat the fixed point argument to get the solution of the thesis $\zeta\in C\left(0,T;W^{s,\frac{6}{5}}(\mathcal{D})\right)$ over the entire time interval $[0,T]$.
+we deduce that, for every $t\in[0,T_{1}]$,
+$\Psi(\zeta_{1})(t)-\Psi(\zeta_{2})(t)=-(-\Delta)^{-\frac{s}{2}}\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(t-r)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\bigl(u_{r}\zeta_{r}\bigr)\,dr.$
+Set
+$\sigma:=\gamma+\frac{s}{2}+\frac{1}{2},\qquad h(r):=(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}(u_{r}\zeta_{r}),$
+and recall that $q:=\frac{p}{\lambda+\delta_{u}}$ and $q^{\prime}=\frac{q}{q-1}$. Using the smoothing estimate for the heat semigroup, we obtain for every $t\in[0,T_{1}]$
+$\begin{split}\|\Psi(\zeta_{1})(t)-\Psi(\zeta_{2})(t)\|_{W^{s,\frac{6}{5}}(\mathcal{D})}&\leq C\int_{0}^{t}\|(-\Delta)^{\sigma}e^{(t-r)\Delta}h(r)\|_{L^{\frac{6}{5}}(\mathcal{D})}\,dr\\ &\leq C\int_{0}^{t}(t-r)^{-\sigma}\|h(r)\|_{L^{\frac{6}{5}}(\mathcal{D})}\,dr.\end{split}$
+Applying Hölder’s inequality in time with exponents $q,q^{\prime}$, we get
+$\|\Psi(\zeta_{1})(t)-\Psi(\zeta_{2})(t)\|_{W^{s,\frac{6}{5}}}$ $\leq C\left(\int_{0}^{t}(t-r)^{-\sigma q^{\prime}}\,dr\right)^{\frac{1}{q^{\prime}}}\left(\int_{0}^{t}\|h(r)\|_{L^{\frac{6}{5}}}^{q}\,dr\right)^{\frac{1}{q}}$
+$\leq C\left(\int_{0}^{T_{1}}\tau^{-\sigma q^{\prime}}\,d\tau\right)^{\frac{1}{q^{\prime}}}\|h\|_{L^{q}(0,T_{1};L^{\frac{6}{5}}(\mathcal{D}))}.$
+The integral in $\tau$ is finite thanks to condition (3.29), i.e. $\sigma q^{\prime}<1$. Hence,
+(3.34) $\|\Psi(\zeta_{1})-\Psi(\zeta_{2})\|_{X}\leq C(\int_{0}^{T_{1}}\tau^{-\sigma q^{\prime}}\,d\tau)^{\frac{1}{q^{\prime}}}\|h\|_{L^{q}(0,T_{1};L^{\frac{6}{5}}(\mathcal{D}))}.$
+Finally, repeating the argument of Step3 in (3.32)-(3.33) we obtain
+$\|h\|_{L^{q}(0,T_{1};L^{\frac{6}{5}}(\mathcal{D}))}\leq C\|u\|_{L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)}\|\zeta\|_{X}.$
+Substituting into (3.34) yields
+(3.35) $\|\Psi(\zeta_{1})-\Psi(\zeta_{2})\|_{X}\leq K(T_{1})\|u\|_{L^{\frac{p}{\lambda+\delta_{u}}}(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))}\|\zeta_{1}-\zeta_{2}\|_{X},$
+where
+$K(T_{1}):=C\left(\int_{0}^{T_{1}}\tau^{-\sigma q^{\prime}}\,d\tau\right)^{\frac{1}{q^{\prime}}}.$
+Since $u\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))$, we have $\|u\|_{L^{\frac{p}{\lambda+\delta_{u}}}(0,T_{1};W^{\frac{1}{2}+\lambda,2})}\to 0$ as $T_{1}\downarrow 0$, and moreover $K(T_{1})<\infty$ for every $T_{1}>0$ by $\sigma q^{\prime}<1$.
+Therefore, choosing $T_{1}>0$ sufficiently small makes the right-hand side of (3.35) strictly less than $\|\zeta_{1}-\zeta_{2}\|_{X}$, so $\Psi$ is a contraction on $X$.
+This yields a unique fixed point $\zeta\in C(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D}))$.
+If $T_{1}<T$, we may iterate the fixed point argument finitely many times to cover $[0,T]$.
+∎
 Remark 3.5.
@@
 We start this section by recalling how the weak Stokes operator $A_{w}$ inherits the property of a bounded $\mathcal{H}^{\infty}$-calculus with $\mathcal{H}^{\infty}$-angle $\varphi^{\infty}_{A_{w}}=0$ from the (strong) Stokes operator $A$.
-This is obtained by following the approach in [25, Section 5].
+This is obtained by following the approach in [27, Section 5].
 From this, it follows the maximal regularity property for the (weak) Stokes operator and the mixed derivative theorem, which are summarised in Proposition 4.2 and Lemma 4.3, respectively.
@@
 ∎
-The first result that we recall is the maximal regularity of the weak Stokes operator, which is a consequence of the $\mathcal{H}^{\infty}$-calculus recalled above, see [24, Section II.4].
+The first result that we recall is the maximal regularity of the weak Stokes operator, which is a consequence of the $\mathcal{H}^{\infty}$-calculus recalled above, see [26, Section II.4].
 Proposition 4.2 (Maximal regularity for the weak-Stokes operator).
@@
     (4.3) $\norm{z}_{\mathbb{E}_{T,p}^{\delta_{u}}}\leq C\left(\norm{g}_{L^{p}(J;H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}))}+\norm{z_{0}}_{V_{p}^{\delta_{u}}}\right).$
-Second, we recall the mixed derivative theorem, see e.g. [24, Corollary 4.5.10].
+Second, we recall the mixed derivative theorem, see e.g. [26, Corollary 4.5.10].
 Lemma 4.3 (Mixed derivative).
@@
 Proof of Theorem 2.8.
-Fix $T>0$, $s\in[0,\frac{1}{2})$ and $p>4$.
-We proceed to verify that parameters satisfying the hypotheses exist.
-First, we choose $\gamma=\frac{1}{4}+\varepsilon_{\gamma}$ with $\varepsilon_{\gamma}>0$ small.
-Recall $a(s,\gamma)=\frac{3}{4}+\frac{s}{2}+\varepsilon_{\gamma}$.
-Since $s<1/2$, we have $a(s,\gamma)<1$ for small $\varepsilon_{\gamma}$.
-We require $\lambda_{\max}>\lambda_{\min}$, i.e., $p(1-a(s,\gamma))>\frac{1}{2}-s$.
-As $\varepsilon_{\gamma}\to 0$, this condition becomes
-$p\left(1-\frac{3}{4}-\frac{s}{2}\right)>\frac{1}{2}-s\iff p\frac{1-2s}{4}>\frac{1-2s}{2}\iff p>2.$
-Since we assumed $p>4>2$, the previous condition is satisfied.
-Next, we choose $\lambda\in(\lambda_{\min},\lambda_{\max})\cap(0,1)$ and define $\overline{\delta}_{\theta},\overline{\delta}_{u}$ as in the statement.
-We must show that the interval for $\delta_{u}$ is non-empty, i.e.,
-$\max\{\delta_{\theta},\tfrac{1}{2}-s\}\leq\delta_{u}<\min\{\overline{\delta}_{u},1-\tfrac{2}{p}\}.$
-We can always choose $\delta_{\theta}$ small enough such that $\delta_{\theta}\leq 1/2-s$.
-Thus, we need to ensure $\frac{1}{2}-s<\min\{\overline{\delta}_{u},1-\frac{2}{p}\}$.
-This is verified, up to choosing $\varepsilon_{\gamma}>0$ small enough, with the same arguments as in Remark 2.9.
-Finally, note that the condition $\delta_{u}<1-2/p$ implies $p>\frac{2}{1-\delta_{u}}$, which ensures the maximal regularity applicability in Theorem 2.5.
+Fix $T>0$, $s\in(0,\tfrac{1}{2})$ and $p>4$. Choose, as in (2.8),
+$\gamma:=\frac{1}{4}+\varepsilon_{\gamma},\qquad 0<\varepsilon_{\gamma}<\min\left\{\frac{1-2s}{4},\frac{(1-2s)(p-4)}{4p}\right\},$
+and set
+$a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2},\qquad\lambda_{\max}:=p(1-a(s,\gamma)),\qquad\lambda_{\min}:=\frac{1}{2}-s.$
+Then $\lambda_{\max}>\lambda_{\min}$, thanks to the same computation as in (2.9).
+Hence the interval $(\lambda_{\min},\lambda_{\max})$ is non-empty.
+Thus, we may choose $\lambda\in(\lambda_{\min},\lambda_{\max})$.
+In particular, as in (2.11), we pick
+$\lambda:=\frac{1}{2}-s+\varepsilon_{\lambda}\qquad 0<\varepsilon_{\lambda}<\min\{2s,\lambda_{\max}-(1-2s)\}.$
+Lastly, as in (2.12), define the thresholds
+$\overline{\delta}_{\theta}:=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\},\qquad\overline{\delta}_{u}:=\min\left\{1-\lambda,\lambda_{\max}-\lambda\right\}.$
+With this choice we have
+$1-\lambda>\frac{1}{2}-s,\qquad\lambda_{\max}-\lambda>\frac{1}{2}-s,\qquad\text{hence}\qquad\overline{\delta}_{u}>\frac{1}{2}-s.$
+In preparation for the application of Sobolev embeddings used below and Theorem 2.5, we now fix $\delta_{u}\in(0,\overline{\delta}_{u})$ and $0<\delta_{\theta}<\overline{\delta}_{\theta}$ satisfying the compatibility conditions (2.15), namely
+$\delta_{u}\geq\max\left\{\delta_{\theta},\frac{1}{2}-s\right\},\qquad\delta_{u}<1-\frac{2}{p}.$
+These constraints are consistent: indeed $p>4$ implies $1-\frac{2}{p}>\frac{1}{2}$ and, by the choice of $\lambda$, we have $\overline{\delta}_{u}>\frac{1}{2}-s$, hence one can pick $\delta_{u}$ in the non-empty interval $\bigl[\frac{1}{2}-s,\min\{\overline{\delta}_{u},1-\frac{2}{p}\}\bigr)$ and then choose any $\delta_{\theta}\in(0,\min\{\overline{\delta}_{\theta},\delta_{u}\})$.
+Finally, note that $\delta_{u}<1-\frac{2}{p}$ is equivalent to $p>\frac{2}{1-\delta_{u}}$, which is the integrability requirement needed to apply maximal regularity in Theorem 2.5.
 Define
 $\alpha_{\theta}=\frac{1}{4}+\frac{\delta_{\theta}}{2},\qquad\beta_{\theta}=\frac{1}{4}-\frac{\delta_{\theta}}{4}.$
-Let $\tilde{\eta}=\tilde{\eta}(p,\delta_{u},T)>0$ and $M=M(p,\delta_{u},T)>0$ be the constants given by Theorem 2.5. Since $\delta_{u}\geq\max\{\delta_{\theta},\frac{1}{2}-s\}$, the embeddings
+Let $\tilde{\eta}=\tilde{\eta}(p,\delta_{u},T)>0$ and $M=M(p,\delta_{u},T)>0$ be the constants given by Theorem 2.5.
+Since $\delta_{u}\geq\max\{\delta_{\theta},\frac{1}{2}-s\}$ thanks to the compatibility conditions (2.15), the embeddings
 $H^{-2\alpha_{\theta}}(\mathcal{D})\hookrightarrow H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}),\quad W^{s,\frac{6}{5}}(\mathcal{D})\hookrightarrow H^{s-1}(\mathcal{D})\hookrightarrow H^{-\frac{1}{2}-\delta_{u}}(\mathcal{D}),$
@@
 […]
-$\tau^{\varepsilon}:=\inf\left\{t>0\ :\ \norm{Z^{\varepsilon}_{t}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}>\frac{\eta}{8\widetilde{M}}\right\}$,
+$\tau^{\varepsilon}:=\inf\left\{t\in[0,T]\ :\ \norm{Z^{\varepsilon}_{t}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}>\frac{\eta}{8\widetilde{M}}\right\}$,
 with the understanding that $\tau^{\varepsilon}=T$ if the previous set is empty. Note that
@@
 To apply the Banach fixed point theorem, we first check that $\Gamma$ maps $\mathbb{B}_{1}$ into itself.
-We can use (2.8) by Corollary 2.3 since its hypothesis, namely $\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $\delta_{u}<\overline{\delta}_{u}$ are satisfied; further, we have the embedding $\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}\hookrightarrow L^{\frac{p}{\lambda+\delta_{u}}}(0,\tau^{\varepsilon};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})),$ which holds for any $0<\lambda\leq 1-\delta_{u}$.
+We can use (2.13) by Corollary 2.3 since its hypothesis, namely $\delta_{\theta}<\overline{\delta}_{\theta}$ and $\delta_{u}<\overline{\delta}_{u}$ are satisfied; further, we have the embedding $\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}\hookrightarrow L^{\frac{p}{\lambda+\delta_{u}}}(0,\tau^{\varepsilon};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})),$ which holds for our fixed $\lambda$ thanks to the assumption $\delta_{u}<1-\lambda$.
 Thus,
@@
 Acknowledgments
+We thank the two anonymous reviewers for their careful reading of the manuscript and for their valuable comments and suggestions, which helped improve the paper.
 G.D.S.
@@
 M.L.
-is supported by the the Italian national inter-university PhD course in Sustainable Development and Climate change.
+is supported by the Italian national inter-university PhD course in Sustainable Development and Climate change.
 M.L.